### v1

# Distinguishing Neighborhood Representations Through Reverse Process of GNNs for Heterophilic Graphs

Abstract
Graph Neural Network (GNN) resembles the diffusion process, leading to the over-smoothing of learned representations when stacking many layers. Hence, the reverse process of message passing can sharpen the node representations by inverting the forward message propagation. The sharpened representations can help us to better distinguish neighboring nodes with different labels, such as in heterophilic graphs. In this work, we apply the design principle of the reverse process to the three variants of the GNNs. Through the experiments on heterophilic graph data, where adjacent nodes need to have different representations for successful classification, we show that the reverse process significantly improves the prediction performance in many cases. Additional analysis reveals that the reverse mechanism can mitigate the over-smoothing over hundreds of layers.
Machine Learning, ICML
[…]
1 Introduction
Graph neural networks (GNN) have emerged as an important tool for learning relational data. Earlier attempts aim to learn the node representations from graphs based on a message-passing mechanism. The message-passing neural network framework shows partial success with the homophilic graphs, where the nodes with the same labels are likely to be connected. When the heterophilic graphs, where node labels significantly differ from those of their neighbors, are considered, the models based on the homophilic assumption often perform worse than the naive neural network architectures without considering the relationship between nodes (Zhu et al., 2020).
Two classes of GNNs have been proposed to overcome the limitation on heterophilic graphs. The first class aims to aggregate the information selectively from the neighbors (Chien et al., 2020; Li et al., 2022). For example, aggregating relevant information from multi-hop neighbors or relevant features from neighbors has been proposed in this direction (Zhu et al., 2020; Liang et al., 2023). On the other hand, the other class of methods highlights the importance of initial node features before aggregation. Adding skip-connection to the initial node features or isolating the original node features for prediction has been proposed in this direction (Zhu et al., 2020; Song et al., 2023).
Figure 1: Visualized node representations over the forward and reverse processes in Minesweeper. Top: our approach with the forward and reverse processes. Bottom: a classical GCN with a forward process only. The original node features are smoothed over the forward process, whereas the features are sharpened over the reverse process. Visualization details are provided in Section 4.1.3.
Aggregation is a fundamental mechanism of all GNN models, and heterophilic GNNs are no exception. With a very clever technique, one can mitigate the propagation of irrelevant information for the first few layers. However, GRAND (Chamberlain et al., 2021) shows that GNN can be seen as a discretization of a heat diffusion equation. The diffusion perspective implies that the node representations will eventually reach the stationary state, with which the nodes cannot be distinguished. Therefore, if we cannot isolate all irrelevant nodes in the aggregation process, which is typically impossible, the GNN will eventually aggregate information from irrelevant nodes as the model propagates over and over.
In this work, we claim not to forcefully correct the diffusive nature of the GNNs. Instead, we propose to use the reverse process of the aggregation. The aggregation process is known to smooth the node representations; hence, its reverse process can sharpen the node representations and make the neighborhood representations more distinguishable. To illustrate our intuition, we showcase our experimental results on the Minesweeper dataset, a well-known heterophilic dataset, in Figure 1. In Minesweeper, a board is a grid-structured graph where each node is initialized with the number of mines in the adjacent nodes $\mathbf{X}^{(0)}$, and the goal is to classify the location of mine $\mathbf{Y}$ correctly. The bottom row visualizes the learned node representations with the GCN (Kipf & Welling, 2017), and the top row visualizes the representations with our approach. With the forward-only method, such as GCN, the learned representations cannot be sharpened enough to distinguish the target nodes from the others, whereas when the reverse process is applied to the initial features, we can obtain a sharpened representation from the backward process.
To this end, we propose the framework of reverse process GNNs utilizing the inversion of forward message-passing layers. Specifically, we provide three variants of reverse process GNN based on three backbone models: 1) GRAND (Chamberlain et al., 2021), 2) GCN (Kipf & Welling, 2017), and 3) GAT (Veličković et al., 2018). For GRAND, we directly use the numerical method to obtain the representations in a backward direction. For GCN and GAT, we adopt the idea of iResNet (Behrmann et al., 2019) to make invertible message-passing layers that can model the reverse process.
The experimental results on heterophilic datasets show that the reverse process improves the prediction performance compared with the forward-only models. Our approach produces sharper representations and successfully stacks deep layers to capture long-range dependencies, which are crucial for performance on heterophilic datasets. The experiments on homophilic datasets confirm that the reverse process does not harm the prediction performance when the aggregation mechanism is sufficient. Our investigation reveals that with the reverse process, one can stack hundreds, even a thousand layers, without suffering over-smoothing. Additional experiments with the low-rate label datasets confirm that the reverse process design is effective for tasks where the long-range interaction between nodes is inevitable.
2 Related Work
Most studies on heterophilic data focus on identifying nodes with similar characteristics even among non-adjacent ones for aggregation. GPR-GNN (Chien et al., 2020) utilizes a trainable generalized PageRank for feature aggregation, learning important neighborhood ranges from the data and emphasizing information within those ranges for aggregation. CPGNN (Zhu et al., 2021) introduces a learnable compatibility matrix to capture the information of non-adjacent homophilic nodes. FAGCN (Bo et al., 2021) introduces a GNN framework with a self-gating mechanism to adaptively use low-frequency and high-frequency signals. FSGNN (Maurya et al., 2022) proposes soft feature selection, wherein it adaptively selects node features with different properties. GloGNN (Li et al., 2022) employs a coefficient matrix that represents node-to-node relationships for aggregation, allowing the aggregation of information from all nodes. GBK-GNN (Du et al., 2022) proposes a bi-kernel graph neural network that separately handles homophilic and heterophilic nodes. It uses a selection gate to predict whether a node is homophilic or heterophilic and obtains features using the corresponding kernel based on the prediction. LRGNN (Liang et al., 2023) uses a low-rank approximation to compute a label relationship matrix, employing it for signed message passing.
However, aggregation still causes global node representations to become similar, known as over-smoothing, leading to performance degradation on heterophilic datasets. To overcome this issue, H2GCN (Zhu et al., 2020) and OGNN (Song et al., 2023) proposes to preserve non-aggregated representation separately. H2GCN learns node representation by separating ego-embedding and neighbor-embedding and employing intermediate representations. OGNN proposes ordering message passing to prevent the mixing of messages from different hops.
On the other hand, several studies tackle the over-smoothing issue. GRAND (Chamberlain et al., 2021) enhances the understanding of over-smoothing from the perspective of the resemblance between GNN structures and the heat diffusion equation. PairNorm (Zhao & Akoglu, 2020) proposes a normalization layer that remains the total pairwise feature distances constant. DropEdge (Rong et al., 2020) randomly removes edges from the graph to cut off messages passing between adjacent nodes. Half-Hop (Azabou et al., 2023) up-sample edges and slow down the message passing. While delaying or limiting over-smoothing can make node representations less smooth, learning difference-enhanced representations between adjacent nodes, which is helpful for heterophilic data, is still challenging.
3 Method
We aim to introduce a framework that can learn distinct node representations between adjacent nodes through reverse diffusion. We first provide the overall framework of our approach and then show three substantiations of the framework for three baseline models.
3.1 Framework
We consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ are a set of nodes and edges respectively, with additional $d$-dimensional node features represented as $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ for all node, and $\mathbf{x}_{i}\in\mathbb{R}^{d}$ denotes the feature of node $i$. It is a well-known fact that a typical GNN layer $f$ tends to learn similar representations between neighboring nodes, leading to the issue of over-smoothing. Chamberlain et al. (2021) highlighted that this is due to the diffusive property inherent in GNN structures.
In contrast to a typical GNN layer, we propose a GNN layer $g$ that performs the opposite role by reversing diffusion and re-concentrating diffused information. Our main idea is to design an inverse function of a message-passing GNN layer $f$, and $g=f^{-1}$ to perform the reversion. Due to the diffusive nature of the GNN layer $f$, its inverse form $g$ is expected to have two properties: 1) $g$ cancels the smoothing effect, producing sharper representations and thus leading to 2) stacking multiple layers of $g$ that do not suffer from the over-smoothing issue.
Formally, with the inverse of multiple message passing layers, our framework predicts node label $\mathbf{Y}$ as follows:
$\hat{\mathbf{Y}}=\phi(f^{(L)}\circ\cdots\circ f^{(1)}(\mathbf{X}^{(0)})\|g^{(1)}\circ\cdots\circ g^{(L)}(\mathbf{X}^{(0)})),$
[…]
In the following sections, we propose a range of methods to develop a reverse diffusion function for GRAND and two variants of GNN with residual connections.
3.2 Reverse Diffusion Based on GRAND
In this section, we suggest a reverse diffusion function based on GRAND. In GRAND, a GNN is interpreted as a discretization of the heat diffusion process. This is modeled by the following heat diffusion equation on the graph:
$\frac{\partial\mathbf{X}(t)}{\partial t}=\big{(}\mathbf{A}(\mathbf{X}(t))-\mathbf{I}\big{)}\mathbf{X}(t),$ (1)
where ${\mathbf{A}}(\mathbf{X})\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ represents the learnable attention matrix. Here, $[A(X)]_{ij}=0$ for any $(i,j)\notin\mathcal{E}$.
Within the framework of diffusion, the time parameter serves as a continuous layer, similar to the concept used in NeuralODEs (Chen et al., 2018). Using Equation 1, GRAND produce node representations at time $T_{D}>0$ by:
[…]
where numerical techniques like the Euler method are used to solve integration. Since Equation 1 models the property of heat reaching equilibrium over time, the node representations obtained through Equation 2 become diffused as time progresses. Conversely, tracing back in time allows us to observe the concentrated form of heat before diffusion. Utilizing Equation 1, past node representations at time $T_{R}<0$ can be calculated as follows:
[…]
Equation 3 reverses the diffusion process, enabling us to obtain sharpened representations.
In experiments, we utilize the GRAND-l model, where the learnable attention matrix remains constant throughout the diffusion process, $\mathbf{A}(\mathbf{X}(t))=\mathbf{A}$, which is known to be parameter-efficient and robust to overfitting. Following the original work, we use the scaled dot product attention to calculate the learnable attention matrix $\mathbf{A}(\mathbf{X})$, which is given as follows:
$[A(X)]_{ij}=\operatorname{softmax}\left(\frac{\left(\mathbf{W}_{\text{K}}\mathbf{x}_{i}\right)^{\top}\mathbf{W}_{\text{Q}}\mathbf{x}_{j}}{d^{\prime}}\right),\;$ (4)
where $\mathbf{W}_{\text{K}}$ and $\mathbf{W}_{\text{Q}}$ are $d^{\prime}\times d$ learnable parameters. When multi-head attention is employed, we use the average of attentions, i.e., $\mathbf{A}(\mathbf{X})=\frac{1}{K}\sum_{k}\mathbf{A}^{(k)}(\mathbf{X})$, where $\mathbf{A}^{(k)}$ is the attention with $k$-th head.
3.3 Reverse Process Based on GNN with residual connections
We have explored the design of reverse process in widely used two message-passing GNN structures: graph convolutional network (GCN) (Kipf & Welling, 2017) and graph attention network (GAT) (Veličković et al., 2018). Both GCN and GAT model a node representation through 1) an aggregation step, where neighborhood representations are combined, and 2) an update step, where the aggregated representations are merged into the target node representation. Let $\hat{\mathbf{A}}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ encode neighborhood structure in a graph, and $\mathbf{W}\in\mathbb{R}^{d\times d}$ be a matrix of learnable parameters. With the application of skip-connections (He et al., 2016), the GCN and GAT layers can be formalized as
$f(\mathbf{X}^{(\ell)})=\mathbf{X}^{(\ell+1)}$ $=\mathbf{X}^{(\ell)}+\sigma(\hat{\mathbf{A}}\mathbf{X}^{(\ell)}\mathbf{W})$
$=\mathbf{X}^{(\ell)}+h(\mathbf{X}^{(\ell)}),\;$
where $\sigma(\cdot)$ represents a non-linear activation function. $\hat{\mathbf{A}}$ is the renormalized adjacency matrix in GCN and a learnable attention matrix in GAT.
According to Behrmann et al. (2019), the inverse of the GNN layer $f^{-1}$ exists if the $\text{Lip}(h)<1$, where $\text{Lip}(h)$ is Lipschitz constant of $h$. In this case, $f^{-1}(\mathbf{X}^{(\ell)})$ can be computed via fixed point iteration as described in Algorithm 1.
Algorithm 1 Inverse of GNN via fixed-point iteration
  Input: output of residual layer $\mathbf{X}^{(\ell)}$, residual block $h$, number of fixed-point iterations $M$
  Output: input of residual layer $\mathbf{X}^{(\ell-1)}$
  $\mathbf{X}\leftarrow\mathbf{X}^{(\ell)}$
  for $m=1,\ldots,M$ do
     $\mathbf{X}\leftarrow\mathbf{X}-h(\mathbf{X})$
  end for
  Return $\mathbf{X}$
If $h$ is a simple feed-forward network, we can constrain its Lipschitz constant by normalizing the weight matrix through the spectral norm. However, computing the spectral norm is time-consuming. Instead, we use an upper bound of $\text{Lip}(h)$ in our case. An upper bound of $\text{Lip}(h)$ is calculated as follows since $\lVert{\mathbf{XY}}\rVert_{F}\leq\lVert{\mathbf{X}}\rVert_{F}\lVert{\mathbf{Y}}\rVert_{F}$:
$\text{Lip}(h)=\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}\mathbf{XW}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}\leq\lVert{\hat{\mathbf{A}}}\rVert_{2}\lVert{\mathbf{W}}\rVert_{F},$ (5)
where $\lVert{\cdot}\rVert_{F}$ denotes Frobenius norm, and $\lVert{\cdot}\rVert_{2}$ spectral norm. In GCN, $\hat{\mathbf{A}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$ where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix $\mathbf{A}$ with added self-loops and $\tilde{\mathbf{D}}$ is the diagonal degree matrix of $\tilde{\mathbf{A}}$. A spectral norm of normalized adjacency matrix $\lVert{\hat{\mathbf{A}}}\rVert_{2}=1$ in GCN structure. Also, $\lVert{\hat{\mathbf{A}}}\rVert_{2}=1$ when $\hat{\mathbf{A}}$ is right-stochastic, which is the case of GAT. Unlike Behrmann et al. (2019), where power iteration is performed to compute the Lipschitz constant, in GCN and GAT, it is sufficient to renormalize the weight matrix through the Frobenius norm of it. The upper bound normalization reduces the time complexity at the expense of the exact Lipschitz constant calculation. In experiments, we find that the Frobenius upper bound can still result in a competitive performance.
When the scaling coefficient $c<1$ is given, $\mathbf{W}$ is normalized to $\frac{c\mathbf{W}}{\lVert{\mathbf{W}}\rVert_{F}}$ if $c<\lVert{\mathbf{W}}\rVert_{F}$, in order to satisfy $\text{Lip}(h)<c$. When multi-head attention with $K$ heads is employed for GAT, we use the average of representations obtained from multiple heads as a node representation. Specifically, parameters of $k$-th head $\mathbf{W}^{(k)}$ is normalized to
$\mathbf{W}^{(k)}\leftarrow\frac{c\mathbf{W}{(k)}}{\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F}},\;\forall k,\;$ (6)
since the Liptschitz upper bound result in $\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F}$. The derivation of Liptschitz upper bound for multi-head attention is provided in Appendix C.
While the time complexity of a GCN mainly depends on the number of the forward propagation $L_{D}$, the complexity of the reverse process depends on the number of fixed point iterations $M$ and of the number of reverse layers $L_{R}$. In our implementation, we run the fixed point iteration until convergence and backpropagate over the iterations. We provide the time and memory complexity analysis in Table 1, and the proof is provided in Appendix A. An analysis on $M$ and the run-time with varying $L_{R}$ in real experiments are provided in Section 4.3.
GCN with residual connection + reverse process
Forward Time $O(L_{D}|\mathcal{E}|d+L_{D}|\mathcal{V}|d^{2})$ $O((L_{D}+M^{2}L_{R})|\mathcal{E}|d+(L_{D}+M^{2}L_{R})|\mathcal{V}|d^{2})$
Forward Memory $O(|\mathcal{E}|+L_{D}|\mathcal{V}|d+d^{2})$ $O(|\mathcal{E}|+(L_{D}+L_{R})|\mathcal{V}|d+d^{2})$
Backward Time $O(L_{D}^{2}|\mathcal{E}|d+L_{D}^{2}|\mathcal{V}|d^{2})$ $O((L_{D}^{2}+M^{2}L_{R}^{3})|\mathcal{E}|d+(L_{D}^{2}+ML_{R}^{2})|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$
Backward Memory $O(|\mathcal{E}|+|\mathcal{V}|d+d^{2})$ $O(|\mathcal{E}|+|\mathcal{V}|d+d^{2})$
Table 1: Space and time complexity of GCN for the forward and backward processes. $L_{D}$ and $L_{R}$ represent the number of forward and reverse steps, respectively, $M$ is the number of fixed point iterations, and $d$ is the dimensionality of the node representation.
4 Experiments
The experimental section focuses on validating two research questions: 1) Can the reverse process produce sharpened representations? 2) Does the reverse process alleviate over-smoothing problems, enabling the construction of deeper layers?
Throughout this section, we denote models with additional reverse layers by ReP (Reverse Process). For example, GCN$+$ReP indicates the GCN backbone with the reverse process. We adopt weight sharing approach of GRAND-l for all experiments.
4.1 Node Classification
In this section, we validate the effectiveness of our framework on node classification. Our primary focus is on assessing performance improvements in heterophilic datasets, while we have also evaluated performance on homophilic datasets.
4.1.1 Datasets
For the node classification task, we utilize a diverse set of datasets to assess our model. For heterophilic data, we explore two Wikipedia graphs, Chameleon and Squirrel, and five additional datasets, Roman-Empire, Amazon-Ratings, Minesweeper, Tolokers, and Questions, introduced by Platonov et al. (2023b). We adopted the filtering process for Chameleon and Squirrel to prevent train-test data leakage as recommended by Platonov et al. (2023b). In the case of homophilic data, our selection includes three citation graphs: Cora, CiteSeer, and PubMed, along with two Amazon co-purchase graphs, Computers and Photo. The statistics of the datasets are summarized in Appendix B.
4.1.2 Experimental Setup and Baselines
For the heterophilic datasets, we adopt the experimental setup from Platonov et al. (2023b), which provides ten random train/validation/test splits. We train a model with cross-entropy loss and report mean accuracy and standard deviation for multi-class classification datasets, including Chameleon, Squirrel, Roman-Empire, and Amazon-Ratings. For binary classification datasets, including Minesweeper, Tolokers, and Questions, binary cross-entropy loss is used, and mean ROC-AUC and standard deviation are reported.
We benchmark several neural architectures as baselines, including classic GNN models like GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and Graph Transformer (GT) (Shi et al., 2020) for more complex attention mechanisms. These baselines are augmented with skip connections and layer normalization. In addition, modifications proposed in Zhu et al. (2020) are made to GAT and GT, resulting in GAT-sep and GT-sep models. For heterophily-specific models, we use 10 models including H2GCN (Zhu et al., 2020), CPGNN (Zhu et al., 2021), GPR-GNN (Chien et al., 2020), FSGNN (Maurya et al., 2022), GloGNN (Li et al., 2022), FAGCN (Bo et al., 2021), GBK-GNN (Du et al., 2022), JacobiConv (Wang & Zhang, 2022), LRGNN (Liang et al., 2023), and OGNN (Song et al., 2023).
For the homophilic datasets, we adopt the experimental setup from He et al. (2021), splitting datasets into 60%/20%/20% train/validation/test sets and using ten random splits for averaging results. We compare our framework against seven baselines: MLP, GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), APPNP (Gasteiger et al., 2018), ChebNet (Defferrard et al., 2016), GPR-GNN (Chien et al., 2020), and BernNet (He et al., 2021).
Validation
For all experiments, we set the number of epochs to 1,000 and apply early stopping when there is no performance improvement for 100 consecutive epochs. For GRAND+ReP, we validate the hyperparameters that maximize the validation metric in the following ranges: learning rate $\in[10^{-5},10^{-1}]$, $T_{D},T_{R}\in[0,10]$, $d\in\{16,32,64,128,256,512\}$, $K\in[1,8]$, $d^{\prime}\in\{4,8,16,32,64,128\}$. With the Euler or Runge-Kutta methods, we search the step size over $[0.5,8]$ and tolerance scale over $[1,20000]$ with the Dormand-Prince method. For GCN+ReP and GAT+ReP, we validate the hyperparameters in the following ranges: learning rate$\in[10^{-5},10^{-1}]$, numbers of forward and reverse layers $L_{D},L_{R}\in\{1,2,4,8,16,32,64,128,256,512\}$, dropout probability $\in[0,0.9]$ with step size of $0.1$, $c\in\{0.1,0.5,0.9,0.999,0.99999\}$, convergence threshold for fixed point iteration $\in\{10^{-4},10^{-5},10^{-6}\}$, $d\in\{128,256,512,1024,2048\}$, and $M\in\{8,16,32,64\}$.
4.1.3 Results
Table 2 shows node classification results on the heterophilic datasets. Results marked with * in Table 2 are obtained from Platonov et al. (2023b) and in Table 3 from He et al. (2021). Applying ReP shows performance improvement for all backbones across most heterophilic datasets, with the most significant and consistent improvement observed in GCN. GCN+ReP achieves state-of-the-art performance in four out of seven datasets and the second-best performance in one. In the datasets where GCN+ReP attained state-of-the-art performance, the number of reverse layers was consistently above 64. This observation shows that deep layers of GNNs mjwith ReP can achieve superior performance without over-smoothing.
Table 3 shows node classification performance on the homophilic datasets. Due to spacing, we only report the results on three datasets. All results are reported in Appendix D. No significant performance changes were observed when ReP applied on homophilic node classification. The results confirm that the sharpened representations do not harm the prediction performance for homophily datasets where the forward aggregation is sufficient.
Squirrel Chameleon Roman- Amazon- Minesweeper Tolokers Questions
empire rating
SAGE* 36.09±1.99 37.77±4.14 85.74±0.67 53.63±0.39 93.51±0.57 82.43±0.44 76.44±0.62
GAT-sep* 35.46±3.10 39.26±2.50 88.75±0.41 52.70±0.62 93.91±0.35 83.78±0.43 76.79±0.71
GT* 36.30±1.98 38.87±3.66 86.51±0.73 51.17±0.66 91.85±0.76 83.23±0.64 77.95±0.68
GT-sep* 36.66±1.63 40.31±3.01 87.32±0.39 52.18±0.80 92.29±0.47 82.52±0.92 78.05±0.93
H2GCN * 35.10±1.15 26.75±3.64 60.11±0.52 36.47±0.23 89.71±0.31 73.35±1.01 63.59±1.46
CPGNN* 30.04±2.03 33.00±3.15 63.96±0.62 39.79±0.77 52.03±5.46 73.36±1.01 65.96±1.95
GPR-GNN* 38.95±1.99 39.93±3.30 64.85±0.27 44.88±0.34 86.24±0.61 72.94±0.97 55.48±0.91
FSGNN* 35.92±1.32 40.61±2.97 79.92±0.56 52.74±0.83 90.08±0.70 82.76±0.61 78.86±0.92
GloGNN* 35.11±1.24 25.90±3.58 59.63±0.69 36.89±0.14 51.08±1.23 73.39±1.17 65.74±1.19
FAGCN* 41.08±2.27 41.90±2.72 65.22±0.56 44.12±0.30 88.17±0.73 77.75±1.05 77.24±1.26
GBK-GNN* 35.51±1.65 39.61±2.60 74.57±0.47 45.98±0.71 90.85±0.58 81.01±0.67 74.47±0.86
JacobiConv* 29.71±1.66 39.00±4.20 71.14±0.42 43.55±0.48 89.66±0.40 68.66±0.65 73.88±1.16
LRGNN 39.51±2.12 41.24±2.95 40.88±1.84 42.23±4.85 52.66±6.40 74.24±1.37 66.41±1.75
OGNN 38.96±2.19 38.04±5.55 80.12±1.22 49.66±1.01 90.21±1.15 81.42±0.65 73.36±1.09
GRAND 35.94±1.64 37.71±4.48 75.19±0.56 49.34±0.72 90.41±0.78 78.38±1.91 76.22±1.06
GRAND+ReP 40.75±2.44 42.14±3.62 77.53±0.62 48.30±0.60 91.42±0.78 80.44±1.64 76.41±1.04
$\Delta$ $+4.81\;(\uparrow)$ $+4.43\;(\uparrow)$ $+2.34\;(\uparrow)$ $-1.04\;(\downarrow)$ $+1.01\;(\uparrow)$ $+2.06\;(\uparrow)$ $+0.19\;(\uparrow)$
GAT* 35.62±2.06 39.21±3.08 80.87±0.30 49.09±0.63 92.01±0.68 83.70±0.47 77.43±1.20
GAT+ReP 39.66±2.00 43.24±4.48 85.87±0.64 52.68±0.27 94.89±0.33 84.52±0.56 76.21±0.74
(32/32)3232(32/32)( 32 / 32 ) $(64/32)$ $(64/4)$ $(16/2)$ $(32/16)$ $(8/1)$ $(64/1)$
$\Delta$ $+4.04\;(\uparrow)$ $+4.03\;(\uparrow)$ $+5.00\;(\uparrow)$ $+3.59\;(\uparrow)$ $+2.11\;(\uparrow)$ $+0.82\;(\uparrow)$ $-1.22\;(\downarrow)$
GCN* 39.47±1.47 40.89±4.12 73.69±0.74 48.70±0.63 89.75±0.52 83.64±0.67 76.09±1.27
GCN+ReP 45.89±1.45 47.57±3.90 86.43±0.74 52.75±0.62 96.05±0.19 86.08±0.84 77.96±0.96
(256/256)256256(256/256)( 256 / 256 ) $(128/256)$ $(256/16)$ $(32/1)$ $(1/256)$ $(128/64)$ $(128/64)$
$\Delta$ $+6.42\;(\uparrow)$ $+6.68\;(\uparrow)$ $+12.74\;(\uparrow)$ $+4.05\;(\uparrow)$ $+6.30\;(\uparrow)$ $+2.44\;(\uparrow)$ $+1.87\;(\uparrow)$
Table 2: Test performance and standard deviation on heterophilic datasets. $\Delta$ indicates the difference with and without ReP. We also report the number of forward and reverse layers below the performance of GCN$+$ReP and GAT$+$ReP. The best and the second-best are bolded and underlined, respectively.
Analysis on the Number of Forward and Reverse Layers
To investigate whether many layers of reverse process improve performance in heterophilic datasets and compare its effect with that of many forward layers, we trained GCN+ReP with varying pairs of steps on two datasets: Chameleon and Minesweeper. Specifically, we vary the number of layers from 1 to 1024 in one direction.
Cora CiteSeer PubMed
MLP* 76.96±0.95 76.58±0.88 85.94±0.22
vanilla-GCN* 87.14±1.01 79.86±0.67 86.74±0.27
vanilla-GAT* 88.03±0.79 80.52±0.71 87.04±0.24
APPNP* 88.14±0.73 80.47±0.74 88.12±0.31
ChevNet* 86.67±0.82 79.11±0.75 87.95±0.28
GPR-GNN* 88.57±0.69 80.12±0.83 88.46±0.33
BernNet* 88.52±0.95 80.09±0.79 88.48±0.41
GRAND 85.53±0.64 74.95±1.37 88.81±0.69
GRAND+ReP 85.73±1.39 75.78±1.48 89.03±0.61
$\Delta$ $+0.20\;(\uparrow)$ $+0.83\;(\uparrow)$ $+0.22\;(\uparrow)$
GAT 87.67±0.84 77.36±1.59 89.66±0.60
GAT+ReP 87.93±1.60 77.06±1.60 89.94±0.61
$(64/32)$ $(32/128)$ (8/8)88(8/8)( 8 / 8 )
$\Delta$ $+0.26(\uparrow)$ $-0.30(\downarrow)$ $+0.28(\uparrow)$
GCN 88.00±1.42 77.15±1.44 89.37±0.52
GCN+ReP 87.63±1.40 77.33±1.65 89.96±0.55
$(32/512)$ $(8/32)$ (32/32)3232(32/32)( 32 / 32 )
$\Delta$ $-0.37(\downarrow)$ $+0.18(\uparrow)$ $+0.59(\uparrow)$
Table 3: Test accuracy and standard deviation on homophilic datasets. $\Delta$ indicates the difference with and without ReP. We also report the number of forward and reverse layers below the performance of GCN$+$ReP and GAT$+$ReP. The best and the second-best are bolded and underlined, respectively.
The prediction performances with varying numbers of layers are reported in Figure 2. In both datasets, the prediction performance keeps increasing as the number of reverse layers increases. These results indicate that the reverse process is capable of deep stacking without suffering from over-smoothing, enabling the models to capture long-range dependencies effectively, which is known to be important, especially in heterophilic graphs. The prediction performance also tends to increase as we increase the number of forward steps up to 1024 in the Chameleon dataset.
Figure 2: Prediction performance with varying numbers of forward and reverse layers. We vary the number of layers in one direction. Due to memory constraints, we restricted the reverse depth used in Minesweeper to 256 or less.
Over-Smoothing Analysis
We evaluate whether the proposed reverse diffusion mitigates the over-smoothing issue. To measure the degree of over-smoothing, we adopt Graph Smoothness Level (GSL) proposed by Zhang et al. (2022) defined as:
$\operatorname{GSL}(\mathbf{X})=\frac{1}{N(N-1)}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V},j\neq i}\frac{\mathbf{x}_{i}\cdot\mathbf{x}_{j}}{\lVert{\mathbf{x}_{i}}\rVert_{F}\lVert{\mathbf{x}_{j}}\rVert_{F}},\;$ (7)
where $\mathbf{x}_{i}$ is the representation of node $i$. The GSL represents the average cosine similarity across all pairs of nodes in the graph. A GSL value closer to one indicates more severe over-smoothing.
Figure 3: Over smoothing levels measured by GSL over the number of layers. Rep (forward) denotes the measured GSL in the forward process of GCN+ReP. We compare the results with GCN of three different numbers of layers: 16, 32, 64.
Figure 3 shows GSL of GCN+ReP and GCN with varying numbers of layers on Squirrel and Chameleon datasets. In both datasets, the GSL of GCN+ReP remains below 0.6 up to 1024 reverse layers, whereas the learned representations from GCN with 32 and 64 layers tend to become similar even after eight layers. GCN with 16 layers shows relatively low GSL values yet still exceeds 0.9 after eight layers. Compared with CGN, GCN+ReP shows relatively less GSL, showing that the reverse process can mitigate the over-smoothing in the forward processes as well.
Qualitative Analysis on Minesweeper Dataset
To validate that the reverse process produces a sharpened representation, we visualize label predictions on the Minesweeper dataset. The Minesweeper dataset is a binary classification task on a grid-structured graph, where the node with a positive label indicates the location of a mine. Each node, unless located on the boundaries, is connected to eight adjacent nodes, including the ones in the diagonal directions. The node feature is initialized with the number of mines in the adjacent nodes, and a one-hot representation of the feature is used as an initial representation of the node for learning.
Based on the representations obtained from the forward and reverse processes, we trained a GCN+ReP model with a single-layer MLP as a prediction head. In Figure 4, we visualize the prediction results of two randomly sampled $7\times 7$ sub-grids from $100\times 100$ grid structure. For each example, we visualize the prediction results with node representations from 1) forward process, 2) reverse process, 3) both directions and 4) true labels displayed from upper left to lower right. Since the prediction head needs concatenated representations for prediction, to visualize the prediction results focused on a forward or reverse representation, we set the other node representation to be zero, e.g., to predict the mine using node representation at layer $\ell>0$, $(\mathbf{X}^{(\ell)}\|\mathbf{0})$ is fed into the prediction head.
[…]
Figure 4: Visualization of node prediction on the minesweeper dataset. We visualize the prediction from 1) forward, 2) reverse, and 3) both representations, along with the ground truth labels.
In both examples, the prediction results from the reverse process appear sharper while those from the forward process tend to be smooth. Additionally, the sharp prediction created by the reverse process significantly contributes to label prediction. Although in some cases, the reverse process can perfectly classify the location of mines, e.g., example 1, the other cases require representations from both directions to classify the mines correctly, e.g., example 2.
Figure 1 shows the changes in predictions over the number of layers. The darker the cell, the higher the predicted probability of a mine being present. As expected, the predictions and representations therein tend to be sharp as the number of reverse layers increases.
4.2 Prediction with Low Label Rate
We compare node classification performance in citation graphs with low label rates, used to mimic the long-range dependencies between nodes. According to Zhao & Akoglu (2020), typical datasets are generally sufficient for shallow architectures, making it difficult to identify the advantages of deep GNNs. As explained by Sun et al. (2020), in scenarios with a low training label rate, more layers are necessary to access distant supervision.
We perform node label classification tasks on the Cora, Citeseer, and Pubmed datasets, with 20 labels per class in the training set. The validation and test sets were consistently set to 500 and 1,000, respectively. Each experiment was conducted across ten random splits of the dataset, with the mean and standard deviation of the results being reported.
We compare node classification accuracy on three citation datasets, Cora, CiteSeer, and PubMed, with baseline methods including GCN and GRAND++ (Thorpe et al., 2022) which is proposed especially for low-labeling rate problems.
Table 4 shows the average accuracy and standard deviation. We find that performance improves when ReP applied. The reverse process is even more effective than GRAND++, a model tailored for the low-rate label prediction task.
Cora CiteSeer PubMed
GRAND++ 72.18±3.10 64.31±1.91 75.94±2.50
GCN 76.50±1.77 61.67±3.30 76.25±2.36
GCN+ReP 79.06±1.54 66.46±1.82 76.92±1.34
Table 4: Prediction accuracy on the citation graphs with 20-label-per-class.
4.3 Run Time Analysis
We first measure how many iterations are required for the fixed point iterations to be converged. Figure 5 shows the difference between consecutive representations over the fixed point iterations with two datasets in terms of mean absolute difference. As shown in the figure, the fixed point method converges after seven iterations in general. In addition, we measure the training time for a single epoch and plot the results in Figure 6. The results show that the training time increases linearly as we increase the number of reverse layers coincided with the complexity analysis in Section 3.3.
Figure 5: The mean absolute difference between two consecutive representations in the fixed-point method.
Figure 6: Average training time for the single epoch. The number of forward layers are shown next to the model.
5 Conclusion
In this work, we propose a reverse process for the message-passing-based graph neural networks. Through extensive empirical analysis, we have found that the reverse process can mitigate over-smoothing issues and allow long-distance nodes to interact with each other. Especially for the heterophilic datasets where the long-range interaction is necessary for a better prediction, the proposed method achieves outstanding results against many baseline models.
Future Work
To ensure invertibility, the Lipschitz constant of the forward process must be restricted to less than one, and the hidden dimension of weight parameters must remain constant. These restrictions limit the representation power and design choices within our framework. Investigating less restrictive invertible forms could lead to performance improvements.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Complexity Analysis
The method for calculating the forward process and the reverse process, when applied to a GCN with residual connections and weight sharing, is as follows:
$\mathbf{X}^{(\ell+1)}=\mathbf{X}^{(\ell)}+\hat{\mathbf{A}}\mathbf{X}^{(\ell)}\mathbf{W},$ (forward process) (8)
$\mathbf{X}^{(\ell-1)}=\mathbf{X}^{(\ell)}+\sum_{k=1}^{K}(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(\ell)}\mathbf{W}^{k},$ (reverse process) (9)
For simplicity, we ignore an activation function.
A.1 Forward Pass of Forward Process
Equation 8 involves three key operations: matrix multiplication between $\mathbf{X}$ and $\mathbf{W}$ with complexity $O(|\mathcal{V}|d^{2})$, matrix multiplication between sparse matrix $\hat{\mathbf{A}}$ and $\mathbf{X}^{(\ell)}\mathbf{W}$ with complexity $O(|\mathcal{E}|d)$, and matrix addition with complexity $O(|\mathcal{V}|d)$. Overall, the time complexity for one-layer is $O(|\mathcal{V}|d^{2}+|\mathcal{E}|d)$. This results in a total time complexity of $O(L_{D}|\mathcal{V}|d^{2}+L_{D}|\mathcal{E}|d)$ over $L_{D}$ layers. Memory cost is calculated as $O(L_{D}|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
A.2 Backward Pass of Reverse Process
The time complexity of the backward pass primarily hinges on the computation of $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$. By applying the chain rule, this can be expressed as:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\frac{\partial\mathbf{X}^{(L_{D})}}{\partial\mathbf{X}^{(L_{D}-1)}}\cdots\frac{\partial\mathbf{X}^{(2)}}{\partial\mathbf{X}^{(1)}}\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}.$ (10)
Given $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}$, $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell-1)}}$ can be derived as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell-1)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\frac{\partial\mathbf{X}^{(\ell)}}{\partial\mathbf{X}^{(\ell-1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\left(\frac{\partial}{\partial\mathbf{X}^{(\ell-1)}}\mathbf{I}\mathbf{X}^{(\ell-1)}+\frac{\partial}{\partial\mathbf{X}^{(\ell-1)}}\hat{\mathbf{A}}\mathbf{X}^{(\ell-1)}\mathbf{W}\right)$
$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}+\hat{\mathbf{A}}^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\mathbf{W}^{\top}.$ (11)
And we can compute $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D}-1)}}$ as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\frac{\partial\mathbf{X}^{(L_{D})}}{\partial\mathbf{X}^{(L_{D}-1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\left(\frac{\partial}{\partial\mathbf{X}^{(L_{D}-1)}}\mathbf{I}\mathbf{X}^{(L_{D}-1)}+\frac{\partial}{\partial\mathbf{X}^{(L_{D}-1)}}\hat{\mathbf{A}}\mathbf{X}^{(L_{D}-1)}\mathbf{W}\right)$
$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}+\hat{\mathbf{A}}^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\mathbf{W}^{\top}.$ (12)
By recursively applying Equation 11 to Equation 12, we can obtain $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(1)}}$ as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\frac{\partial\mathbf{X}^{(L_{D})}}{\partial\mathbf{X}^{(L_{D}-1)}}\cdots\frac{\partial\mathbf{X}^{(2)}}{\partial\mathbf{X}^{(1)}}=\sum_{\ell=0}^{L_{D}-1}\binom{L_{D}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\right)({\mathbf{W}^{\top}})^{\ell}.$ (13)
Finally, $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ can be derived as the multiplication of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(1)}}$ and $\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}$:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\bigg{(}\sum_{\ell=0}^{L_{D}-1}\binom{L_{D}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\right)({\mathbf{W}^{\top}})^{\ell}\bigg{)}\left(\frac{\partial}{\partial\mathbf{W}}\hat{\mathbf{A}}\mathbf{X}^{(0)}\mathbf{W}\right)$
$=\left(\hat{\mathbf{A}}\mathbf{X}^{(0)}\right)^{\top}\bigg{(}\sum_{\ell=0}^{L_{D}-1}\binom{L_{D}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\right)({\mathbf{W}^{\top}})^{\ell}\bigg{)}.$ (14)
Note that $\hat{\mathbf{A}}$ is the sparse matrix, so matrix multiplication can be performed in $O(|\mathcal{E}|d)$. So, the time complexity of the term in the summation is $O(|\mathcal{V}|d^{2}+\ell d^{3}+\ell|\mathcal{E}|d)$. Thus, the time complexity for computing $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ is $O(L_{D}|\mathcal{V}|d^{2}+L_{D}^{2}d^{3}+L_{D}^{2}|\mathcal{E}|d)$, with a memory cost of $O(|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
A.3 Forward Pass of Reverse Process
To calculate Equation 9, the initial step involves computing $(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(\ell)}\mathbf{W}^{k}$. The time complexity for this part is $O(k|\mathcal{E}|d+|\mathcal{V}|d^{2}+kd^{3})$. Summing over $k$ from 0 to $K$, the overall complexity becomes $O(K^{2}|\mathcal{E}|d+K|\mathcal{V}|d^{2}+K^{2}d^{3})$. With $L_{R}$ representing the number of reverse process layers, the total time complexity is calculated as $O(K^{2}L_{R}|\mathcal{E}|d+KL_{R}|\mathcal{V}|d^{2}+K^{2}L_{R}d^{3})$.
A.4 Backward Pass of Reverse Process
The time complexity of the backward pass primarily hinges on the computation of $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$. By applying the chain rule, this can be expressed as:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}\cdots\frac{\partial\mathbf{X}^{(-2)}}{\partial\mathbf{X}^{(-1)}}\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}.$ (15)
Given $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}$, $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell+1)}}$ can be derived as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell+1)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\frac{\partial\mathbf{X}^{(\ell)}}{\partial\mathbf{X}^{(\ell+1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\frac{\partial}{\partial\mathbf{X}^{(\ell+1)}}\left(\mathbf{X}^{(\ell+1)}+\sum_{k=1}^{K}(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(\ell+1)}\mathbf{W}^{k}\right)$
$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}+\sum_{k=1}^{K}(-1)^{k}\left(\hat{\mathbf{A}}^{k}\right)^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\left(\mathbf{W}^{k}\right)^{\top}.$ (16)
And we can compute $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R}+1)}}$ as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R}+1)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial}{\partial\mathbf{X}^{(-L_{R}+1)}}\left(\mathbf{X}^{(-L_{R}+1)}+\sum_{k=1}^{K}(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(-L_{R}+1)}\mathbf{W}^{k}\right)$
$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{k=1}^{K}(-1)^{k}\left(\hat{\mathbf{A}}^{k}\right)^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\left(\mathbf{W}^{k}\right)^{\top}.$ (17)
By recursively applying Equation 16 to Equation 17, we can obtain $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-1)}}$ as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}\cdots\frac{\partial\mathbf{X}^{(-2)}}{\partial\mathbf{X}^{(-1)}}$
$=$ $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{\ell=2}^{L_{R}-1}\binom{L_{R}-1}{\ell}\sum_{(k_{1},k_{2},\cdots,k_{\ell-1})}(-1)^{\sum_{i=1}^{\ell-1}k_{i}}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}.$ (18)
In addition, $\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$ can be obtained as follows:
$\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$ $=\frac{\partial}{\partial\mathbf{W}}\left(\mathbf{X}^{(0)}+\sum_{k_{0}=1}^{K}(-1)^{k_{0}}\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)}\mathbf{W}^{k_{0}}\right)$
$=\sum_{k_{0}=1}^{K}(-1)^{k_{0}}k_{0}\frac{\partial}{\partial\mathbf{W}^{k_{0}}}(\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)}\mathbf{W}^{k_{0}})\mathbf{W}^{k_{0}-1}\frac{\partial\mathbf{W}}{\partial\mathbf{W}}.$ (19)
Finally, $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ can be derived as the multiplication of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-1)}}$ and $\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\sum_{k_{0}=1}^{K}\Bigg{[}(-1)^{k_{0}}k_{0}(\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)})^{\top}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)\mathbf{W}^{k_{0}-1}$
$+\sum_{\ell=2}^{L_{R}-1}\binom{L_{R}-1}{\ell}\sum_{(k_{1},k_{2},\cdots,k_{\ell-1})}(-1)^{\sum_{i=1}^{\ell-1}k_{i}}k_{0}(\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)})^{\top}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}\mathbf{W}^{k_{0}-1}\Bigg{]}.$ (20)
By the same procedure in the forward process, the time complexity of the first term in brackets is $O(k_{0}(|\mathcal{E}|d+d^{3})+|\mathcal{V}|d^{2})$. In the case of the second term, there are up to $2\ell K$ possible outcomes for $\sum_{i=1}^{\ell}k_{i}$. So the time complexity is $O(L_{R}+K\ell|\mathcal{V}|d^{2}+\ell^{2}K^{2}(d^{3}+|\mathcal{E}|d))$. So the overall time complexity of reverse process is $O((L_{D}^{2}+M^{2}L_{R}^{3})|\mathcal{E}|d+(L_{D}^{2}+ML_{R}^{2})|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$. The memory complexity is $O(|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
Appendix B Dataset Statistics
Table 5 presents the dataset statistics utilized in experiments. There are two forms of homophily: edge homophily (Abu-El-Haija et al., 2019; Zhu et al., 2020) and adjusted homophily (Platonov et al., 2023a). Edge homophily denotes the proportion of edges connecting nodes with the same label, formally expressed as:
$h_{\text{edge }}=\frac{\left.\mid(u,v)\in E:y_{u}=y_{v}\right\}\mid}{|E|},$
where $E$ is the set of edges, and $y_{n}$ is the label of node $n$. However, edge homophily is acknowledged to be meaningless in graphs with imbalanced labels. To address this issue, adjusted homophily is introduced. Formally, adjusted homophily is defined as:
$h_{adj}=\frac{h_{\text{edge }}-\sum_{k=1}^{C}D_{k}^{2}/(2|E|)^{2}}{1-\sum_{k=1}^{C}D_{k}^{2}/(2|E|)^{2}},$
where $D_{k}$ is the total degree of nodes of class $k$, and $C$ is the number of classes. We employed seven heterophilic datasets characterized by low adjusted homophily and five commonly used homophilic datasets exhibiting high edge homophily and adjusted homophily.
Dataset # nodes # edges # classes avg degree edge homophily adjusted homophily
Squirrel-filtered 2,223 46,998 5 42.28 0.21 0.01
Chameleon-filtered 890 8,854 5 19.90 0.24 0.03
roman-empire 22,662 32,927 18 2.91 0.05 -0.05
amazon-ratings 24,492 93,050 5 7.60 0.38 0.14
minesweeper 10,000 39,402 2 7.88 0.68 0.01
tolokers 11,758 519,000 2 88.28 0.59 0.09
questions 48,921 153,540 2 6.28 0.84 0.02
Cora 2,708 5,278 7 3.90 0.81 0.77
CiteSeer 3,327 4,552 6 2.74 0.74 0.67
PubMed 19,717 44,324 3 4.50 0.80 0.69
Computers 13,752 245,861 10 35.76 0.78 0.68
Photo 7,650 119,081 8 31.13 0.83 0.79
Table 5: Statistics of the dataset utilized in the experiments.
Appendix C Multi-head Attention for GAT
GAT calculates an attention matrix as follows:
$\hat{A}_{ij}=\frac{\exp\left(\operatorname{LeakyReLU}\left(\mathbf{a}^{\top}\left[\mathbf{W}^{\top}\mathbf{X}_{i}\|\mathbf{W}^{\top}\mathbf{X}_{j}\right]\right)\right)}{\sum_{k\in\mathcal{N}_{i}}\exp\left(\operatorname{LeakyReLU}\left(\mathbf{a}^{\top}\left[\mathbf{W}^{\top}\mathbf{X}_{i}\|\mathbf{W}^{\top}\mathbf{X}_{j}\right]\right)\right)},\;$ (21)
where $\mathbf{a}\in\mathbb{R}^{2d}$ is a learnable parameter. Our framework also adopts averaging when using multi-head attention and remains hidden dimension constant to ensure invertibility, resulting in:
$h(\mathbf{X}^{(\ell)})=\sigma(\frac{1}{K}\sum_{k=1}^{K}\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}\mathbf{W}^{(k)}).\;$ (22)
In this case, the upper bound of the Lipschitz constant of $h$ is calculated as follows:
$\text{Lip}(h)$ $=\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\frac{1}{K}\sum_{k=1}^{K}\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}\mathbf{W}^{(k)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}$ (23)
$\leq\frac{1}{K}\sum_{k=1}^{K}\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}\lVert{\mathbf{W}^{(k)}}\rVert_{F}$ (24)
$\leq\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F},\;$ (25)
since $\lVert{\sum\mathbf{X}}\rVert_{F}\leq\sum\lVert{\mathbf{X}}\rVert_{F}$.
Appendix D Full Results of Homophily Datasets
We provide the experimental results of all homophilic datasets in Table 6.
Cora CiteSeer PubMed Computers Photo
MLP* 76.96±0.95 76.58±0.88 85.94±0.22 82.85±0.38 84.72±0.34
vanilla-GCN* 87.14±1.01 79.86±0.67 86.74±0.27 83.32±0.33 88.26±0.73
vanilla-GAT* 88.03±0.79 80.52±0.71 87.04±0.24 83.32±0.39 90.94±0.68
APPNP* 88.14±0.73 80.47±0.74 88.12±0.31 85.32±0.37 88.51±0.31
ChevNet* 86.67±0.82 79.11±0.75 87.95±0.28 87.54±0.43 93.77±0.32
GPR-GNN* 88.57±0.69 80.12±0.83 88.46±0.33 86.85±0.25 93.85±0.28
BernNet* 88.52±0.95 80.09±0.79 88.48±0.41 87.64±0.44 93.63±0.35
GRAND 85.53±0.64 74.95±1.37 88.81±0.69 90.28±0.47 94.01±0.73
GRAND+ReP 85.73±1.39 75.78±1.48 89.03±0.61 89.51±0.78 94.48±0.61
$\Delta$ $+0.20\;(\uparrow)$ $+0.83\;(\uparrow)$ $+0.22\;(\uparrow)$ $-0.77\;(\downarrow)$ $+0.47\;(\uparrow)$
GAT 87.67±0.84 77.36±1.59 89.66±0.60 92.15±0.30 95.86±0.58
GAT+ReP 87.93±1.60 77.06±1.60 89.94±0.61 91.03±0.62 95.44±0.71
$(64/32)$ $(32/128)$ (8/8)88(8/8)( 8 / 8 ) $(16/8)$ $(32/8)$
$\Delta$ $+0.26(\uparrow)$ $-0.30(\downarrow)$ $+0.28(\uparrow)$ $-1.12(\downarrow)$ $-0.42(\downarrow)$
GCN 88.00±1.42 77.15±1.44 89.37±0.52 91.87±0.57 95.35±0.47
GCN+ReP 87.63±1.40 77.33±1.65 89.96±0.55 90.92±0.52 95.50±0.63
$(32/512)$ $(8/32)$ ($2/32)( 3$$2/64)( 3$2/128)32128(32/128)( 32 / 128 )
$\Delta$ $-0.37(\downarrow)$ $+0.18(\uparrow)$ $+0.59(\uparrow)$ $-0.95(\downarrow)$ $+0.15(\uparrow)$
Table 6: Test accuracy and standard deviation on homophilic datasets. $\Delta$ indicates the difference with and without ReP. We also report the number of forward and reverse layers below the performance of GCN$+$ReP and GAT$+$ReP. The best and the second-best are bolded and underlined, respectively.

### v2

@@
-Distinguishing Neighborhood Representations Through Reverse Process of GNNs for Heterophilic Graphs
+Mitigating Oversmoothing Through Reverse Process of GNNs for Heterophilic Graphs
 Abstract
 Graph Neural Network (GNN) resembles the diffusion process, leading to the over-smoothing of learned representations when stacking many layers.
-Hence, the reverse process of message passing can sharpen the node representations by inverting the forward message propagation.
-The sharpened representations can help us to better distinguish neighboring nodes with different labels, such as in heterophilic graphs.
+Hence, the reverse process of message passing can produce the distinguishable node representations by inverting the forward message propagation.
+The distinguishable representations can help us to better classify neighboring nodes with different labels, such as in heterophilic graphs.
 In this work, we apply the design principle of the reverse process to the three variants of the GNNs.
@@
 Additional analysis reveals that the reverse mechanism can mitigate the over-smoothing over hundreds of layers.
+Our code is available at https://github.com/ml-postech/reverse-gnn.
 Machine Learning, ICML
-[…]
 1 Introduction
-Graph neural networks (GNN) have emerged as an important tool for learning relational data.
+Graph neural networks (GNNs) have emerged as an important tool for learning relational data.
 Earlier attempts aim to learn the node representations from graphs based on a message-passing mechanism.
 The message-passing neural network framework shows partial success with the homophilic graphs, where the nodes with the same labels are likely to be connected.
-When the heterophilic graphs, where node labels significantly differ from those of their neighbors, are considered, the models based on the homophilic assumption often perform worse than the naive neural network architectures without considering the relationship between nodes (Zhu et al., 2020).
-Two classes of GNNs have been proposed to overcome the limitation on heterophilic graphs.
-The first class aims to aggregate the information selectively from the neighbors (Chien et al., 2020; Li et al., 2022).
-For example, aggregating relevant information from multi-hop neighbors or relevant features from neighbors has been proposed in this direction (Zhu et al., 2020; Liang et al., 2023).
-On the other hand, the other class of methods highlights the importance of initial node features before aggregation.
-Adding skip-connection to the initial node features or isolating the original node features for prediction has been proposed in this direction (Zhu et al., 2020; Song et al., 2023).
+When the heterophilic graphs, where node labels significantly differ from those of their neighbors, are considered, the models based on the homophilic assumption (McPherson et al., 2001) often perform worse than the naive neural network architectures without considering the relationship between nodes (Zhu et al., 2020).
 Figure 1: Visualized node representations over the forward and reverse processes in Minesweeper.
@@
 Bottom: a classical GCN with a forward process only.
-The original node features are smoothed over the forward process, whereas the features are sharpened over the reverse process.
+The original node features are smoothed over the forward process, whereas the features are more distinguishable over the reverse process.
 Visualization details are provided in Section 4.1.3.
-Aggregation is a fundamental mechanism of all GNN models, and heterophilic GNNs are no exception.
-With a very clever technique, one can mitigate the propagation of irrelevant information for the first few layers.
-However, GRAND (Chamberlain et al., 2021) shows that GNN can be seen as a discretization of a heat diffusion equation.
-The diffusion perspective implies that the node representations will eventually reach the stationary state, with which the nodes cannot be distinguished.
-Therefore, if we cannot isolate all irrelevant nodes in the aggregation process, which is typically impossible, the GNN will eventually aggregate information from irrelevant nodes as the model propagates over and over.
+To learn the node representations of heterophilic graphs, a GNN needs to capture long-range interactions between nodes, leading to the stacking of multiple message-passing layers (Li et al., 2022; Rusch et al., 2023).
+However, many studies empirically and theoretically identify that GNN tends to smooth the node representations over the layers, and eventually, the learned representations are likely to be similar, known as the over-smoothing issue (Chen et al., 2020; Rusch et al., 2023).
+Furthermore, GRAND (Chamberlain et al., 2021) shows that GNN can be seen as a discretization of a heat diffusion equation.
+The diffusion perspective implies that learning distinguishable node representation with deep GNN is challenging since heat only diffuses to reach equilibrium, where node representation becomes indistinguishable.
 In this work, we claim not to forcefully correct the diffusive nature of the GNNs.
 Instead, we propose to use the reverse process of the aggregation.
-The aggregation process is known to smooth the node representations; hence, its reverse process can sharpen the node representations and make the neighborhood representations more distinguishable.
-To illustrate our intuition, we showcase our experimental results on the Minesweeper dataset, a well-known heterophilic dataset, in Figure 1.
+The aggregation process is known to make the node representations similar; hence, its reverse process can make the neighborhood representations more distinguishable.
+Revisiting the diffusion perspective, applying the reverse process means learning the states in the past, which are further away from equilibrium and more distinguishable.
+To illustrate our intuition, we showcase our experimental results on the Minesweeper dataset, a well-known heterophilic dataset (Platonov et al., 2023b), in Figure 1.
 In Minesweeper, a board is a grid-structured graph where each node is initialized with the number of mines in the adjacent nodes $\mathbf{X}^{(0)}$, and the goal is to classify the location of mine $\mathbf{Y}$ correctly.
-The bottom row visualizes the learned node representations with the GCN (Kipf & Welling, 2017), and the top row visualizes the representations with our approach.
-With the forward-only method, such as GCN, the learned representations cannot be sharpened enough to distinguish the target nodes from the others, whereas when the reverse process is applied to the initial features, we can obtain a sharpened representation from the backward process.
+The top row visualizes the learned node representations with our approach, and the bottom row visualizes the representations with the GCN (Kipf & Welling, 2017).
+With the forward-only method, such as GCN, the learned representations often fail to obtain distinguishable representations for classification.
+However, when the reverse process is applied to the initial features, we can obtain a distinguishable representation from the backward process.
 To this end, we propose the framework of reverse process GNNs utilizing the inversion of forward message-passing layers.
@@
 The experimental results on heterophilic datasets show that the reverse process improves the prediction performance compared with the forward-only models.
-Our approach produces sharper representations and successfully stacks deep layers to capture long-range dependencies, which are crucial for performance on heterophilic datasets.
+Our investigation reveals that the reverse process produces distinguishable representations and enables the stacking of hundreds, even a thousand layers, mitigating over-smoothing.
+Successfully stacking deep layers allows for the capture of long-range dependencies, which are crucial for performance on heterophilic datasets.
 The experiments on homophilic datasets confirm that the reverse process does not harm the prediction performance when the aggregation mechanism is sufficient.
-Our investigation reveals that with the reverse process, one can stack hundreds, even a thousand layers, without suffering over-smoothing.
-Additional experiments with the low-rate label datasets confirm that the reverse process design is effective for tasks where the long-range interaction between nodes is inevitable.
 2 Related Work
@@
 CPGNN (Zhu et al., 2021) introduces a learnable compatibility matrix to capture the information of non-adjacent homophilic nodes.
-FAGCN (Bo et al., 2021) introduces a GNN framework with a self-gating mechanism to adaptively use low-frequency and high-frequency signals.
-FSGNN (Maurya et al., 2022) proposes soft feature selection, wherein it adaptively selects node features with different properties.
+FSGNN (Maurya et al., 2022) proposes soft feature selection, wherein it adaptively selects neighbors to aggregate with different hop distances.
 GloGNN (Li et al., 2022) employs a coefficient matrix that represents node-to-node relationships for aggregation, allowing the aggregation of information from all nodes.
@@
 However, aggregation still causes global node representations to become similar, known as over-smoothing, leading to performance degradation on heterophilic datasets.
-To overcome this issue, H2GCN (Zhu et al., 2020) and OGNN (Song et al., 2023) proposes to preserve non-aggregated representation separately.
+To overcome this issue, H2GCN (Zhu et al., 2020) and Ordered GNN (Song et al., 2023) proposes to preserve non-aggregated representation separately.
 H2GCN learns node representation by separating ego-embedding and neighbor-embedding and employing intermediate representations.
-OGNN proposes ordering message passing to prevent the mixing of messages from different hops.
+Ordered GNN proposes ordering message passing to prevent the mixing of messages from different hops.
+Several studies propose methods to adaptively learn appropriate filters that can handle various graph structures.
+FAGCN (Bo et al., 2021) introduces a GNN framework with a self-gating mechanism to adaptively use low-frequency and high-frequency signals.
+JacobiConv (Wang & Zhang, 2022) uses Jacobi bases for spectral filter, whose orthogonality and flexibility enable adaptation to a wide range of graph signal densities.
+ACM-GCN (Luan et al., 2022) utilizes a filterbank which combines low-pass and high-pass filters, and adaptively considers node-wise local information.
 On the other hand, several studies tackle the over-smoothing issue.
@@
 In contrast to a typical GNN layer, we propose a GNN layer $g$ that performs the opposite role by reversing diffusion and re-concentrating diffused information.
-Our main idea is to design an inverse function of a message-passing GNN layer $f$, and $g=f^{-1}$ to perform the reversion.
-Due to the diffusive nature of the GNN layer $f$, its inverse form $g$ is expected to have two properties: 1) $g$ cancels the smoothing effect, producing sharper representations and thus leading to 2) stacking multiple layers of $g$ that do not suffer from the over-smoothing issue.
+Our main idea is to design an inverse function of a message-passing GNN layer $f$, with $g=f^{-1}$ to perform the reversion.
+Due to the diffusive nature of the GNN layer $f$, its inverse form $g$ is expected to have two properties: 1) $g$ cancels the smoothing effect, producing distinguishable representations that mitigate the over-smoothing issue and thus 2) leading to stacking multiple layers of $g$.
 Formally, with the inverse of multiple message passing layers, our framework predicts node label $\mathbf{Y}$ as follows:
@@
 Within the framework of diffusion, the time parameter serves as a continuous layer, similar to the concept used in NeuralODEs (Chen et al., 2018).
-Using Equation 1, GRAND produce node representations at time $T_{D}>0$ by:
+Using Equation 1, GRAND produce node representations at time $T_{F}>0$ by:
 […]
@@
 Conversely, tracing back in time allows us to observe the concentrated form of heat before diffusion.
-Utilizing Equation 1, past node representations at time $T_{R}<0$ can be calculated as follows:
+Utilizing Equation 1, node representations at a past time $T_{R}<0$ can be calculated as follows:
 […]
-Equation 3 reverses the diffusion process, enabling us to obtain sharpened representations.
+Equation 3 reverses the diffusion process, enabling us to obtain distinguishable representations.
 In experiments, we utilize the GRAND-l model, where the learnable attention matrix remains constant throughout the diffusion process, $\mathbf{A}(\mathbf{X}(t))=\mathbf{A}$, which is known to be parameter-efficient and robust to overfitting.
@@
 We have explored the design of reverse process in widely used two message-passing GNN structures: graph convolutional network (GCN) (Kipf & Welling, 2017) and graph attention network (GAT) (Veličković et al., 2018).
-Both GCN and GAT model a node representation through 1) an aggregation step, where neighborhood representations are combined, and 2) an update step, where the aggregated representations are merged into the target node representation.
+Both GCN and GAT model a node representation through an aggregation step, where neighborhood representations are combined, and an update step, where the aggregated representations are merged into the target node representation.
 Let $\hat{\mathbf{A}}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ encode neighborhood structure in a graph, and $\mathbf{W}\in\mathbb{R}^{d\times d}$ be a matrix of learnable parameters.
 With the application of skip-connections (He et al., 2016), the GCN and GAT layers can be formalized as
-$f(\mathbf{X}^{(\ell)})=\mathbf{X}^{(\ell+1)}$ $=\mathbf{X}^{(\ell)}+\sigma(\hat{\mathbf{A}}\mathbf{X}^{(\ell)}\mathbf{W})$
-$=\mathbf{X}^{(\ell)}+h(\mathbf{X}^{(\ell)}),\;$
-where $\sigma(\cdot)$ represents a non-linear activation function. $\hat{\mathbf{A}}$ is the renormalized adjacency matrix in GCN and a learnable attention matrix in GAT.
+$f(\mathbf{X}^{(\ell)})=\mathbf{X}^{(\ell+1)}$ $=\mathbf{X}^{(\ell)}+\sigma(\hat{\mathbf{A}}\mathbf{X}^{(\ell)}\mathbf{W})$ (5)
+$=\mathbf{X}^{(\ell)}+h(\mathbf{X}^{(\ell)}),\;$ (6)
+where $\sigma(\cdot)$ represents a non-linear activation function.
+$\hat{\mathbf{A}}$ is the renormalized adjacency matrix in GCN and a learnable attention matrix in GAT.
+We note that when using multi-head attention, we take averaging approach as in GRAND to keep invertibility.
 According to Behrmann et al.
-(2019), the inverse of the GNN layer $f^{-1}$ exists if the $\text{Lip}(h)<1$, where $\text{Lip}(h)$ is Lipschitz constant of $h$.
-In this case, $f^{-1}(\mathbf{X}^{(\ell)})$ can be computed via fixed point iteration as described in Algorithm 1.
+(2019), the inverse of the GNN layer $g=f^{-1}$ exists if the $\text{Lip}(h)<1$, where $\text{Lip}(h)$ is Lipschitz constant of $h$.
+With the contractive nonlinear activations like ReLU, ELU, and tanh, $\text{Lip}(h)<1$ is satisfied if
+$\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}\mathbf{XW}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}<1\;,$ (7)
+where $\lVert{\cdot}\rVert_{F}$ denotes Frobenius norm.
+When the condition is guaranteed, $g(\mathbf{X}^{(\ell)})$ can be computed via fixed point iteration as described in Algorithm 1, resulting in $\mathbf{X}^{(\ell-1)}$.
+To ensure the invertibility of $f$ throughout the entire training procedure, we enforce the weight matrix $\mathbf{W}$ to satisfy the condition.
+Since it is difficult to optimize the weight matrix while satisfying the condition, we normalize the weight matrix after each gradient descent step.
+Specifically, given that the left side of Equation 7 is upper bounded by
+$\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}\mathbf{XW}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}\leq\lVert{\hat{\mathbf{A}}}\rVert_{2}\lVert{\mathbf{W}}\rVert_{F}\;,$ (8)
+where $\lVert{\cdot}\rVert_{2}$ denotes spectral norm, we use the upper bound to normalize the weight matrix.
+Note that the spectral norm of $\hat{\mathbf{A}}$ is straightforward for the two models that we considered as baselines: GCN and GAT.
+In GCN, $\hat{\mathbf{A}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$ where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix $\mathbf{A}$ with added self-loops and $\tilde{\mathbf{D}}$ is the diagonal degree matrix of $\tilde{\mathbf{A}}$.
+A spectral norm of renormalized adjacency matrix $\lVert{\hat{\mathbf{A}}}\rVert_{2}=1$.
+In GAT, $\lVert{\hat{\mathbf{A}}}\rVert_{2}=1$ since $\hat{\mathbf{A}}$ is right-stochastic.
+Therefore, normalizing the weight matrix through its Frobenius norm is sufficient to guarantee the condition in Equation 7.
+The upper bound normalization reduces the time complexity at the expense of the exact supremum calculation.
+In experiments, we find that the Frobenius upper bound can still result in a competitive performance.
 Algorithm 1 Inverse of GNN via fixed-point iteration
-  Input: output of residual layer $\mathbf{X}^{(\ell)}$, residual block $h$, number of fixed-point iterations $M$
+  Input: output of residual layer $\mathbf{X}^{(\ell)}$, residual block $h$, the number of fixed-point iterations $M$
   Output: input of residual layer $\mathbf{X}^{(\ell-1)}$
@@
   Return $\mathbf{X}$
-If $h$ is a simple feed-forward network, we can constrain its Lipschitz constant by normalizing the weight matrix through the spectral norm.
-However, computing the spectral norm is time-consuming.
-Instead, we use an upper bound of $\text{Lip}(h)$ in our case.
-An upper bound of $\text{Lip}(h)$ is calculated as follows since $\lVert{\mathbf{XY}}\rVert_{F}\leq\lVert{\mathbf{X}}\rVert_{F}\lVert{\mathbf{Y}}\rVert_{F}$:
-$\text{Lip}(h)=\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}\mathbf{XW}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}\leq\lVert{\hat{\mathbf{A}}}\rVert_{2}\lVert{\mathbf{W}}\rVert_{F},$ (5)
-where $\lVert{\cdot}\rVert_{F}$ denotes Frobenius norm, and $\lVert{\cdot}\rVert_{2}$ spectral norm.
-In GCN, $\hat{\mathbf{A}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$ where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix $\mathbf{A}$ with added self-loops and $\tilde{\mathbf{D}}$ is the diagonal degree matrix of $\tilde{\mathbf{A}}$.
-A spectral norm of normalized adjacency matrix $\lVert{\hat{\mathbf{A}}}\rVert_{2}=1$ in GCN structure.
-Also, $\lVert{\hat{\mathbf{A}}}\rVert_{2}=1$ when $\hat{\mathbf{A}}$ is right-stochastic, which is the case of GAT.
-Unlike Behrmann et al.
-(2019), where power iteration is performed to compute the Lipschitz constant, in GCN and GAT, it is sufficient to renormalize the weight matrix through the Frobenius norm of it.
-The upper bound normalization reduces the time complexity at the expense of the exact Lipschitz constant calculation.
-In experiments, we find that the Frobenius upper bound can still result in a competitive performance.
 When the scaling coefficient $c<1$ is given, $\mathbf{W}$ is normalized to $\frac{c\mathbf{W}}{\lVert{\mathbf{W}}\rVert_{F}}$ if $c<\lVert{\mathbf{W}}\rVert_{F}$, in order to satisfy $\text{Lip}(h)<c$.
-When multi-head attention with $K$ heads is employed for GAT, we use the average of representations obtained from multiple heads as a node representation.
-Specifically, parameters of $k$-th head $\mathbf{W}^{(k)}$ is normalized to
-$\mathbf{W}^{(k)}\leftarrow\frac{c\mathbf{W}{(k)}}{\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F}},\;\forall k,\;$ (6)
-since the Liptschitz upper bound result in $\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F}$. The derivation of Liptschitz upper bound for multi-head attention is provided in Appendix C.
-While the time complexity of a GCN mainly depends on the number of the forward propagation $L_{D}$, the complexity of the reverse process depends on the number of fixed point iterations $M$ and of the number of reverse layers $L_{R}$.
+When multi-head attention with $K$ heads is employed for GAT, parameters of $k$-th head $\mathbf{W}^{(k)}$ is normalized to $\frac{c\mathbf{W}^{(k)}}{\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F}}$ for all $k$, since the upper bound result in $\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F}$.
+The derivation of the upper bound for multi-head attention is provided in Appendix C.
+Since residual block $h$ is an operator on a Banach space, and we constraint the $\text{Lip}(h)<1$, the convergence of Algorithm 1 is guaranteed by the Banach fixed point theorem (Behrmann et al., 2019).
+Inversion error in practice are reported in Section 4.2.
+While the time complexity of a GCN mainly depends on the number of the forward layers $L_{F}$, the complexity of the reverse process depends on the number of fixed point iterations $M$ and of the number of reverse layers $L_{R}$.
 In our implementation, we run the fixed point iteration until convergence and backpropagate over the iterations.
@@
 GCN with Residual Connection + Reverse Process
-Forward Time $O(L_{D}|\mathcal{E}|d+L_{D}|\mathcal{V}|d^{2})$ $O((L_{D}+M^{2}L_{R})|\mathcal{E}|d+(L_{D}+M^{2}L_{R})|\mathcal{V}|d^{2})$
-Forward Memory $O(|\mathcal{E}|+L_{D}|\mathcal{V}|d+d^{2})$ $O(|\mathcal{E}|+(L_{D}+L_{R})|\mathcal{V}|d+d^{2})$
-Backward Time $O(L_{D}^{2}|\mathcal{E}|d+L_{D}^{2}|\mathcal{V}|d^{2})$ $O((L_{D}^{2}+M^{2}L_{R}^{3})|\mathcal{E}|d+(L_{D}^{2}+ML_{R}^{2})|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$
-Backward Memory $O(|\mathcal{E}|+|\mathcal{V}|d+d^{2})$ $O(|\mathcal{E}|+|\mathcal{V}|d+d^{2})$
-Table 1: Space and time complexity of GCN for the forward and backward processes.
-$L_{D}$ and $L_{R}$ represent the number of forward and reverse steps, respectively, $M$ is the number of fixed point iterations, and $d$ is the dimensionality of the node representation.
+Forward Time $O(L_{F}|\mathcal{E}|d+L_{F}|\mathcal{V}|d^{2})$ $+\;O(M^{2}L_{R}|\mathcal{E}|d+ML_{R}|\mathcal{V}|d^{2}+M^{2}L_{R}d^{3})$
+Forward Memory $O(|\mathcal{E}|+L_{F}|\mathcal{V}|d+d^{2})$ $+\;O(L_{R}|\mathcal{V}|d)$
+Backward Time $O(L_{F}^{2}|\mathcal{E}|d+L_{F}|\mathcal{V}|d^{2}+L_{F}^{2}d^{3})$ $+\;O(M^{2}L_{R}^{3}|\mathcal{E}|d+ML_{R}^{2}|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$
+Backward Memory $O(|\mathcal{E}|+|\mathcal{V}|d+d^{2})$
+Table 1: Space and time complexity of GCN for the forward and reverse processes.
+We show an additional complexity when using reverse process, for simplicity.
+✗ denotes there are no additional complexity.
+$L_{F}$ and $L_{R}$ represent the number of forward and reverse layers, respectively, $M$ is the number of fixed point iterations, and $d$ is the dimensionality of the node representation.
 4 Experiments
-The experimental section focuses on validating two research questions: 1) Can the reverse process produce sharpened representations?
+The experimental section focuses on validating two research questions: 1) Can the reverse process produce distinguishable representations?
 2) Does the reverse process alleviate over-smoothing problems, enabling the construction of deeper layers?
@@
 For example, GCN$+$ReP indicates the GCN backbone with the reverse process.
-We adopt weight sharing approach of GRAND-l for all experiments.
+We adopt weight sharing approach of GRAND-l for all experiments using ReP.
 4.1 Node Classification
@@
 (2020) are made to GAT and GT, resulting in GAT-sep and GT-sep models.
-For heterophily-specific models, we use 10 models including H2GCN (Zhu et al., 2020), CPGNN (Zhu et al., 2021), GPR-GNN (Chien et al., 2020), FSGNN (Maurya et al., 2022), GloGNN (Li et al., 2022), FAGCN (Bo et al., 2021), GBK-GNN (Du et al., 2022), JacobiConv (Wang & Zhang, 2022), LRGNN (Liang et al., 2023), and OGNN (Song et al., 2023).
+For heterophily-specific models, we use 10 models including H2GCN (Zhu et al., 2020), CPGNN (Zhu et al., 2021), GPR-GNN (Chien et al., 2020), FSGNN (Maurya et al., 2022), GloGNN (Li et al., 2022), FAGCN (Bo et al., 2021), GBK-GNN (Du et al., 2022), JacobiConv (Wang & Zhang, 2022), LRGNN (Liang et al., 2023), Ordered GNN (Song et al., 2023), ACM-GCN (Luan et al., 2022), and Dir-GNN (Rossi et al., 2023).
 For the homophilic datasets, we adopt the experimental setup from He et al.
@@
 For all experiments, we set the number of epochs to 1,000 and apply early stopping when there is no performance improvement for 100 consecutive epochs.
-For GRAND+ReP, we validate the hyperparameters that maximize the validation metric in the following ranges: learning rate $\in[10^{-5},10^{-1}]$, $T_{D},T_{R}\in[0,10]$, $d\in\{16,32,64,128,256,512\}$, $K\in[1,8]$, $d^{\prime}\in\{4,8,16,32,64,128\}$.
+For GRAND+ReP, we validate the hyperparameters that maximize the validation metric in the following ranges: learning rate $\in[10^{-5},10^{-1}]$, $T_{F},T_{R}\in[0,10]$, $d\in\{16,32,64,128,256,512\}$, $K\in[1,8]$, $d^{\prime}\in\{4,8,16,32,64,128\}$.
 With the Euler or Runge-Kutta methods, we search the step size over $[0.5,8]$ and tolerance scale over $[1,20000]$ with the Dormand-Prince method.
-For GCN+ReP and GAT+ReP, we validate the hyperparameters in the following ranges: learning rate$\in[10^{-5},10^{-1}]$, numbers of forward and reverse layers $L_{D},L_{R}\in\{1,2,4,8,16,32,64,128,256,512\}$, dropout probability $\in[0,0.9]$ with step size of $0.1$, $c\in\{0.1,0.5,0.9,0.999,0.99999\}$, convergence threshold for fixed point iteration $\in\{10^{-4},10^{-5},10^{-6}\}$, $d\in\{128,256,512,1024,2048\}$, and $M\in\{8,16,32,64\}$.
+For GCN+ReP and GAT+ReP, we validate the hyperparameters in the following ranges: learning rate$\in[10^{-5},10^{-1}]$, the number of forward and reverse layers $L_{F},L_{R}\in\{1,2,4,8,16,32,64,128,256,512\}$, dropout probability $\in[0,0.9]$ with step size of $0.1$, $c\in\{0.1,0.5,0.9,0.999,0.99999\}$, convergence threshold for fixed point iteration $\in\{10^{-4},10^{-5},10^{-6}\}$, $d\in\{128,256,512,1024,2048\}$, and $M\in\{8,16,32,64\}$.
+We fix the non-linear activation function to ReLU.
 4.1.3 Results
-Table 2 shows node classification results on the heterophilic datasets.
-Results marked with * in Table 2 are obtained from Platonov et al.
-(2023b) and in Table 3 from He et al.
-(2021).
-Applying ReP shows performance improvement for all backbones across most heterophilic datasets, with the most significant and consistent improvement observed in GCN.
-GCN+ReP achieves state-of-the-art performance in four out of seven datasets and the second-best performance in one.
-In the datasets where GCN+ReP attained state-of-the-art performance, the number of reverse layers was consistently above 64.
-This observation shows that deep layers of GNNs mjwith ReP can achieve superior performance without over-smoothing.
-Table 3 shows node classification performance on the homophilic datasets.
-Due to spacing, we only report the results on three datasets.
-All results are reported in Appendix D.
-No significant performance changes were observed when ReP applied on homophilic node classification.
-The results confirm that the sharpened representations do not harm the prediction performance for homophily datasets where the forward aggregation is sufficient.
 Squirrel Chameleon Roman- Amazon- Minesweeper Tolokers Questions
-empire rating
+empire ratings
 SAGE* 36.09±1.99 37.77±4.14 85.74±0.67 53.63±0.39 93.51±0.57 82.43±0.44 76.44±0.62
@@
 LRGNN 39.51±2.12 41.24±2.95 40.88±1.84 42.23±4.85 52.66±6.40 74.24±1.37 66.41±1.75
-OGNN 38.96±2.19 38.04±5.55 80.12±1.22 49.66±1.01 90.21±1.15 81.42±0.65 73.36±1.09
+Ordered GNN 38.96±2.19 38.04±5.55 80.12±1.22 49.66±1.01 90.21±1.15 81.42±0.65 73.36±1.09
+ACM-GCN 33.07±3.03 31.78±3.35 69.66±0.62 32.26±2.06 90.53±0.56 79.18±0.77 62.50±4.05
+Dir-GNN 40.39±1.11 41.26±2.00 91.23±0.32 44.88±0.84 91.35±0.65 81.78±0.83 76.30±0.99
 GRAND 35.94±1.64 37.71±4.48 75.19±0.56 49.34±0.72 90.41±0.78 78.38±1.91 76.22±1.06
@@
 The best and the second-best are bolded and underlined, respectively.
+Table 2 shows node classification results on the heterophilic datasets.
+Results marked with * and in Table 2 are obtained from Platonov et al.
+(2023b) and Rossi et al.
+(2023), and * in Table 3 from He et al.
+(2021).
+Applying ReP shows performance improvement for all backbones across most heterophilic datasets, with the most significant and consistent improvement observed in GCN.
+GCN+ReP achieves state-of-the-art performance in four out of seven datasets and the second-best performance in one.
+In the datasets where GCN+ReP attained state-of-the-art performance, the number of reverse layers was consistently above 64.
+This observation shows that deep layers of GNNs with ReP can achieve superior performance without over-smoothing.
+Table 3 shows node classification performance on the homophilic datasets.
+Due to spacing, we only report the results on three datasets.
+All results are reported in Appendix D.
+No significant performance changes were observed when ReP applied on homophilic node classification.
+The results confirm that the distinguishable representations do not harm the prediction performance for homophily datasets where the forward aggregation is sufficient.
 Analysis on the Number of Forward and Reverse Layers
@@
 In both datasets, the prediction performance keeps increasing as the number of reverse layers increases.
-These results indicate that the reverse process is capable of deep stacking without suffering from over-smoothing, enabling the models to capture long-range dependencies effectively, which is known to be important, especially in heterophilic graphs.
+These results indicate that the reverse process is capable of deep stacking to mitigate over-smoothing.
+This enables the models to capture long-range dependencies effectively, which is known to be important, especially in heterophilic graphs.
 The prediction performance also tends to increase as we increase the number of forward steps up to 1024 in the Chameleon dataset.
-Figure 2: Prediction performance with varying numbers of forward and reverse layers.
-We vary the number of layers in one direction.
+Figure 2: Prediction performance with varying the number of forward and reverse layers.
+We vary the number of layers (depth) in one direction.
 Due to memory constraints, we restricted the reverse depth used in Minesweeper to 256 or less.
 Over-Smoothing Analysis
-We evaluate whether the proposed reverse diffusion mitigates the over-smoothing issue.
+We evaluate whether the proposed reverse process mitigates the over-smoothing issue.
 To measure the degree of over-smoothing, we adopt Graph Smoothness Level (GSL) proposed by Zhang et al.
 (2022) defined as:
-$\operatorname{GSL}(\mathbf{X})=\frac{1}{N(N-1)}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V},j\neq i}\frac{\mathbf{x}_{i}\cdot\mathbf{x}_{j}}{\lVert{\mathbf{x}_{i}}\rVert_{F}\lVert{\mathbf{x}_{j}}\rVert_{F}},\;$ (7)
+$\operatorname{GSL}(\mathbf{X})=\frac{1}{|\mathcal{V}|(|\mathcal{V}|-1)}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V},j\neq i}\frac{\mathbf{x}_{i}\cdot\mathbf{x}_{j}}{\lVert{\mathbf{x}_{i}}\rVert_{F}\lVert{\mathbf{x}_{j}}\rVert_{F}},\;$ (9)
 where $\mathbf{x}_{i}$ is the representation of node $i$.
@@
 A GSL value closer to one indicates more severe over-smoothing.
-Figure 3: Over smoothing levels measured by GSL over the number of layers.
+Figure 3: Over-smoothing levels measured by GSL over the number of layers (depth).
 ReP (forward) denotes the measured GSL in the forward process of GCN+ReP.
-We compare the results with GCN of three different numbers of layers: 16, 32, 64.
+We compare the results with GCN of three different depths: 16, 32, 64.
 Figure 3 shows GSL of GCN+ReP and GCN with varying numbers of layers on Squirrel and Chameleon datasets.
@@
 Qualitative Analysis on Minesweeper Dataset
-To validate that the reverse process produces a sharpened representation, we visualize label predictions on the Minesweeper dataset.
+To validate that the reverse process produces a distinguishable representation, we visualize label predictions on the Minesweeper dataset.
 The Minesweeper dataset is a binary classification task on a grid-structured graph, where the node with a positive label indicates the location of a mine.
@@
 In Figure 4, we visualize the prediction results of two randomly sampled $7\times 7$ sub-grids from $100\times 100$ grid structure.
-For each example, we visualize the prediction results with node representations from 1) forward process, 2) reverse process, 3) both directions and 4) true labels displayed from upper left to lower right.
+For each example, we visualize the prediction results with node representations from 1) forward process, 2) reverse process, and 3) both directions, as well as 4) the true labels, displayed from upper left to lower right.
+In the visualization of the true labels, black cells indicate the presence of mine, and white cells indicate its absence.
+In the visualization of the prediction results, the darker the cell, the higher the predicted probability of a mine being present.
 Since the prediction head needs concatenated representations for prediction, to visualize the prediction results focused on a forward or reverse representation, we set the other node representation to be zero, e.g., to predict the mine using node representation at layer $\ell>0$, $(\mathbf{X}^{(\ell)}\|\mathbf{0})$ is fed into the prediction head.
-[…]
-[…]
-Figure 4: Visualization of node prediction on the minesweeper dataset. We visualize the prediction from 1) forward, 2) reverse, and 3) both representations, along with the ground truth labels.
-In both examples, the prediction results from the reverse process appear sharper while those from the forward process tend to be smooth.
-Additionally, the sharp prediction created by the reverse process significantly contributes to label prediction.
+(a) Example 1
+(b) Example 2
+Figure 4: Visualization of node prediction on the Minesweeper dataset. We visualize the prediction from 1) forward, 2) reverse, and 3) both representations, along with 4) the ground truth labels.
+In both examples, the prediction results from the reverse process appear distinguishable, while those from the forward process tend to be smooth.
+Additionally, the distinguishable prediction created by the reverse process significantly contributes to label prediction.
 Although in some cases, the reverse process can perfectly classify the location of mines, e.g., Example 1, the other cases require representations from both directions to classify the mines correctly, e.g., Example 2.
-Figure 1 shows the changes in predictions over the number of layers.
-The darker the cell, the higher the predicted probability of a mine being present.
-As expected, the predictions and representations therein tend to be sharp as the number of reverse layers increases.
-4.2 Prediction with Low Label Rate
-We compare node classification performance in citation graphs with low label rates, used to mimic the long-range dependencies between nodes.
-According to Zhao & Akoglu (2020), typical datasets are generally sufficient for shallow architectures, making it difficult to identify the advantages of deep GNNs.
-As explained by Sun et al.
-(2020), in scenarios with a low training label rate, more layers are necessary to access distant supervision.
-We perform node label classification tasks on the Cora, Citeseer, and Pubmed datasets, with 20 labels per class in the training set.
-The validation and test sets were consistently set to 500 and 1,000, respectively.
-Each experiment was conducted across ten random splits of the dataset, with the mean and standard deviation of the results being reported.
-We compare node classification accuracy on three citation datasets, Cora, CiteSeer, and PubMed, with baseline methods including GCN and GRAND++ (Thorpe et al., 2022) which is proposed especially for low-labeling rate problems.
-Table 4 shows the average accuracy and standard deviation.
-We find that performance improves when ReP applied.
-The reverse process is even more effective than GRAND++, a model tailored for the low-rate label prediction task.
-Cora CiteSeer PubMed
-GRAND++ 72.18±3.10 64.31±1.91 75.94±2.50
-GCN 76.50±1.77 61.67±3.30 76.25±2.36
-GCN+ReP 79.06±1.54 66.46±1.82 76.92±1.34
-Table 4: Prediction accuracy on the citation graphs with 20-label-per-class.
+Figure 1 shows the changes in predictions over the number of layers on the $10\times 10$ sampled sub-grids.
+We follow the same visualization procedure with Figure 4.
+As expected, the predictions and representations tend to be more distinguishable as the number of reverse layers increases.
+4.2 Inversion Error
+Although the invertibility of Algorithm 1 is guaranteed in theory, the inversion may not be achieved due to numerical errors.
+To verify the fixed point method, we conduct the experiment to restore the inputs from the outputs of GCN and GAT at a depth of $64$.
+We set the scaling coefficient $c$ to $0.99999$ and the number of fixed point iterations to eight, which is challenging due to the large coefficient (note that the Lipschitz of 1 is non-invertible) and the small number of iterations.
+Table 4 shows the mean absolute error between the original inputs and restored input data.
+As the results show, the inversion error is negligible in practice.
+GCN+ReP GAT+ReP
+Squirrel $3.36\times 10^{-5}$ $3.28\times 10^{-5}$
+Chameleon $2.23\times 10^{-5}$ $2.41\times 10^{-5}$
+Roman-empire $2.79\times 10^{-5}$ $3.56\times 10^{-5}$
+Amazon-ratings $4.31\times 10^{-5}$ $3.46\times 10^{-5}$
+Table 4: Mean absolute error between the original inputs and restored input data by Algorithm 1.
 4.3 Run Time Analysis
@@
 Future Work
-To ensure invertibility, the Lipschitz constant of the forward process must be restricted to less than one, and the hidden dimension of weight parameters must remain constant.
-These restrictions limit the representation power and design choices within our framework.
+To ensure invertibility, the Lipschitz constant of the forward process must be restricted, and the hidden dimension of weight parameters must remain constant.
+These restrictions limit the representation power and design choices.
 Investigating less restrictive invertible forms could lead to performance improvements.
@@
 There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+Acknowledgements
+This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)) and supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00337955 and RS-2023-00217286).
 Appendix A Complexity Analysis
-The method for calculating the forward process and the reverse process, when applied to a GCN with residual connections and weight sharing, is as follows:
-$\mathbf{X}^{(\ell+1)}=\mathbf{X}^{(\ell)}+\hat{\mathbf{A}}\mathbf{X}^{(\ell)}\mathbf{W},$ (forward process) (8)
-$\mathbf{X}^{(\ell-1)}=\mathbf{X}^{(\ell)}+\sum_{k=1}^{K}(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(\ell)}\mathbf{W}^{k},$ (reverse process) (9)
-For simplicity, we ignore an activation function.
+The forward and reverse processes of a GCN with residual connections and weight sharing are as follows:
+[…]
+[…]
+where $M$ is the number of fixed point iterations. For simplicity, we ignore an activation function.
 A.1 Forward Pass of Forward Process
 Equation 10 involves three key operations: matrix multiplication between $\mathbf{X}$ and $\mathbf{W}$ with complexity $O(|\mathcal{V}|d^{2})$, matrix multiplication between sparse matrix $\hat{\mathbf{A}}$ and $\mathbf{X}^{(\ell)}\mathbf{W}$ with complexity $O(|\mathcal{E}|d)$, and matrix addition with complexity $O(|\mathcal{V}|d)$.
-Overall, the time complexity for one-layer is $O(|\mathcal{V}|d^{2}+|\mathcal{E}|d)$.
-This results in a total time complexity of $O(L_{D}|\mathcal{V}|d^{2}+L_{D}|\mathcal{E}|d)$ over $L_{D}$ layers.
-Memory cost is calculated as $O(L_{D}|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
-A.2 Backward Pass of Reverse Process
-The time complexity of the backward pass primarily hinges on the computation of $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$. By applying the chain rule, this can be expressed as:
-$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\frac{\partial\mathbf{X}^{(L_{D})}}{\partial\mathbf{X}^{(L_{D}-1)}}\cdots\frac{\partial\mathbf{X}^{(2)}}{\partial\mathbf{X}^{(1)}}\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}.$ (10)
-Given $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}$, $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell-1)}}$ can be derived as follows:
-$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell-1)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\frac{\partial\mathbf{X}^{(\ell)}}{\partial\mathbf{X}^{(\ell-1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\left(\frac{\partial}{\partial\mathbf{X}^{(\ell-1)}}\mathbf{I}\mathbf{X}^{(\ell-1)}+\frac{\partial}{\partial\mathbf{X}^{(\ell-1)}}\hat{\mathbf{A}}\mathbf{X}^{(\ell-1)}\mathbf{W}\right)$
-$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}+\hat{\mathbf{A}}^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\mathbf{W}^{\top}.$ (11)
-And we can compute $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D}-1)}}$ as follows:
-$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\frac{\partial\mathbf{X}^{(L_{D})}}{\partial\mathbf{X}^{(L_{D}-1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\left(\frac{\partial}{\partial\mathbf{X}^{(L_{D}-1)}}\mathbf{I}\mathbf{X}^{(L_{D}-1)}+\frac{\partial}{\partial\mathbf{X}^{(L_{D}-1)}}\hat{\mathbf{A}}\mathbf{X}^{(L_{D}-1)}\mathbf{W}\right)$
-$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}+\hat{\mathbf{A}}^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\mathbf{W}^{\top}.$ (12)
-By recursively applying Equation 11 to Equation 12, we can obtain $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(1)}}$ as follows:
-$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\frac{\partial\mathbf{X}^{(L_{D})}}{\partial\mathbf{X}^{(L_{D}-1)}}\cdots\frac{\partial\mathbf{X}^{(2)}}{\partial\mathbf{X}^{(1)}}=\sum_{\ell=0}^{L_{D}-1}\binom{L_{D}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\right)({\mathbf{W}^{\top}})^{\ell}.$ (13)
-Finally, $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ can be derived as the multiplication of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(1)}}$ and $\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}$:
-$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\bigg{(}\sum_{\ell=0}^{L_{D}-1}\binom{L_{D}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\right)({\mathbf{W}^{\top}})^{\ell}\bigg{)}\left(\frac{\partial}{\partial\mathbf{W}}\hat{\mathbf{A}}\mathbf{X}^{(0)}\mathbf{W}\right)$
-$=\left(\hat{\mathbf{A}}\mathbf{X}^{(0)}\right)^{\top}\bigg{(}\sum_{\ell=0}^{L_{D}-1}\binom{L_{D}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\right)({\mathbf{W}^{\top}})^{\ell}\bigg{)}.$ (14)
-Note that $\hat{\mathbf{A}}$ is the sparse matrix, so matrix multiplication can be performed in $O(|\mathcal{E}|d)$.
-So, the time complexity of the term in the summation is $O(|\mathcal{V}|d^{2}+\ell d^{3}+\ell|\mathcal{E}|d)$.
-Thus, the time complexity for computing $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ is $O(L_{D}|\mathcal{V}|d^{2}+L_{D}^{2}d^{3}+L_{D}^{2}|\mathcal{E}|d)$, with a memory cost of $O(|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
+Overall, the time complexity for a single layer is $O(|\mathcal{V}|d^{2}+|\mathcal{E}|d)$.
+Therefore, the total time complexity over $L_{F}$ forward layers is $O(L_{F}|\mathcal{V}|d^{2}+L_{F}|\mathcal{E}|d)$.
+Memory cost is calculated as $O(L_{F}|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
+A.2 Backward Pass of Forward Process
+The time complexity of the backward pass is primarily determined by the computation cost of $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$.
+By applying the chain rule, the gradient can be expressed as:
+$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\frac{\partial\mathbf{X}^{(L_{F})}}{\partial\mathbf{X}^{(L_{F}-1)}}\cdots\frac{\partial\mathbf{X}^{(2)}}{\partial\mathbf{X}^{(1)}}\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}.$ (12)
+By sequentially multiplying $\frac{\partial\mathbf{X}^{(n)}}{\partial\mathbf{X}^{(n-1)}}$ for $n=L_{F},\cdots,2$ on the right side of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}$, we derive:
+$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\frac{\partial\mathbf{X}^{(L_{F})}}{\partial\mathbf{X}^{(L_{F}-1)}}\cdots\frac{\partial\mathbf{X}^{(2)}}{\partial\mathbf{X}^{(1)}}=\sum_{\ell=0}^{L_{F}-1}\binom{L_{F}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\right)({\mathbf{W}^{\top}})^{\ell}\;.$ (13)
+To demonstrate Equation 13, we illustrate part of the sequential multiplication process.
+Multiplying $\frac{\partial\mathbf{X}^{(L_{F})}}{\partial\mathbf{X}^{(L_{F}-1)}}$ to the right side of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}$, we get:
+$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\frac{\partial\mathbf{X}^{(L_{F})}}{\partial\mathbf{X}^{(L_{F}-1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\left(\frac{\partial}{\partial\mathbf{X}^{(L_{F}-1)}}\mathbf{I}\mathbf{X}^{(L_{F}-1)}+\frac{\partial}{\partial\mathbf{X}^{(L_{F}-1)}}\hat{\mathbf{A}}\mathbf{X}^{(L_{F}-1)}\mathbf{W}\right)$
+$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}+\hat{\mathbf{A}}^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\mathbf{W}^{\top}.$ (14)
+Next, we multiply $\frac{\partial\mathbf{X}^{(L_{F}-1)}}{\partial\mathbf{X}^{(L_{F}-2)}}$ on the right side of Equation 14, yielding:
+$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\frac{\partial\mathbf{X}^{(L_{F})}}{\partial\mathbf{X}^{(L_{F}-1)}}\frac{\partial\mathbf{X}^{(L_{F}-1)}}{\partial\mathbf{X}^{(L_{F}-2)}}$ $=\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}+\hat{\mathbf{A}}^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\mathbf{W}^{\top}\right)\left(\frac{\partial}{\partial\mathbf{X}^{(L_{F}-2)}}\mathbf{I}\mathbf{X}^{(L_{F}-2)}+\frac{\partial}{\partial\mathbf{X}^{(L_{F}-2)}}\hat{\mathbf{A}}\mathbf{X}^{(L_{F}-2)}\mathbf{W}\right)$
+$=\sum_{\ell=0}^{2}\binom{2}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\right)({\mathbf{W}^{\top}})^{\ell}\;.$ (15)
+Repeating the process above, we can derive Equation 13.
+Finally, $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ is derived by multiplying Equation 13 with $\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}$:
+$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(1)}}\right)\left(\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}\right)$
+$=\bigg{(}\sum_{\ell=0}^{L_{F}-1}\binom{L_{F}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\right)({\mathbf{W}^{\top}})^{\ell}\bigg{)}\left(\frac{\partial}{\partial\mathbf{W}}\hat{\mathbf{A}}\mathbf{X}^{(0)}\mathbf{W}\right)$
+$=\left(\hat{\mathbf{A}}\mathbf{X}^{(0)}\right)^{\top}\bigg{(}\sum_{\ell=0}^{L_{F}-1}\binom{L_{F}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\right)({\mathbf{W}^{\top}})^{\ell}\bigg{)}.$ (16)
+The time complexity of the term in the summation is $O(L_{F}+\ell|\mathcal{E}|d+|\mathcal{V}|d^{2}+\ell d^{3})$.
+Therefore, the overall time complexity for computing $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ is $O(L_{F}^{2}d^{3}+L_{F}|\mathcal{V}|d^{2}+L_{F}^{2}|\mathcal{E}|d)$, with a memory cost of $O(|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
 A.3 Forward Pass of Reverse Process
-To calculate Equation 9, the initial step involves computing $(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(\ell)}\mathbf{W}^{k}$.
-The time complexity for this part is $O(k|\mathcal{E}|d+|\mathcal{V}|d^{2}+kd^{3})$.
-Summing over $k$ from 0 to $K$, the overall complexity becomes $O(K^{2}|\mathcal{E}|d+K|\mathcal{V}|d^{2}+K^{2}d^{3})$.
-With $L_{R}$ representing the number of reverse process layers, the total time complexity is calculated as $O(K^{2}L_{R}|\mathcal{E}|d+KL_{R}|\mathcal{V}|d^{2}+K^{2}L_{R}d^{3})$.
+To calculate Equation 11, the initial step involves computing $(-1)^{m}\hat{\mathbf{A}}^{m}\mathbf{X}^{(\ell)}\mathbf{W}^{m}$.
+The time complexity for this part is $O(m|\mathcal{E}|d+|\mathcal{V}|d^{2}+md^{3})$.
+Summing over $m$ from 0 to $M$, the overall complexity becomes $O(M^{2}|\mathcal{E}|d+M|\mathcal{V}|d^{2}+M^{2}d^{3})$.
+With $L_{R}$ representing the number of reverse process layers, additional time complexity becomes $O(M^{2}L_{R}|\mathcal{E}|d+ML_{R}|\mathcal{V}|d^{2}+M^{2}L_{R}d^{3})$ comparing to forward pass without reverse process.
 A.4 Backward Pass of Reverse Process
-The time complexity of the backward pass primarily hinges on the computation of $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$. By applying the chain rule, this can be expressed as:
+The time complexity of the backward pass is primarily determined by the computation cost of $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$.
+By applying the chain rule, the gradient can be expressed as:
 $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}\cdots\frac{\partial\mathbf{X}^{(-2)}}{\partial\mathbf{X}^{(-1)}}\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}.$ (17)
-Given $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}$, $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell+1)}}$ can be derived as follows:
-$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell+1)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\frac{\partial\mathbf{X}^{(\ell)}}{\partial\mathbf{X}^{(\ell+1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\frac{\partial}{\partial\mathbf{X}^{(\ell+1)}}\left(\mathbf{X}^{(\ell+1)}+\sum_{k=1}^{K}(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(\ell+1)}\mathbf{W}^{k}\right)$
-$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}+\sum_{k=1}^{K}(-1)^{k}\left(\hat{\mathbf{A}}^{k}\right)^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\left(\mathbf{W}^{k}\right)^{\top}.$ (16)
-And we can compute $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R}+1)}}$ as follows:
-$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R}+1)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial}{\partial\mathbf{X}^{(-L_{R}+1)}}\left(\mathbf{X}^{(-L_{R}+1)}+\sum_{k=1}^{K}(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(-L_{R}+1)}\mathbf{W}^{k}\right)$
-$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{k=1}^{K}(-1)^{k}\left(\hat{\mathbf{A}}^{k}\right)^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\left(\mathbf{W}^{k}\right)^{\top}.$ (17)
-By recursively applying Equation 16 to Equation 17, we can obtain $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-1)}}$ as follows:
+By sequentially multiplying $\frac{\partial\mathbf{X}^{(-n)}}{\partial\mathbf{X}^{(-n+1)}}$ for $n=L_{R},\cdots,2$ to the right side of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}$, we derive:
 $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}\cdots\frac{\partial\mathbf{X}^{(-2)}}{\partial\mathbf{X}^{(-1)}}$
-$=$ $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{\ell=2}^{L_{R}-1}\binom{L_{R}-1}{\ell}\sum_{(k_{1},k_{2},\cdots,k_{\ell-1})}(-1)^{\sum_{i=1}^{\ell-1}k_{i}}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}.$ (18)
-In addition, $\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$ can be obtained as follows:
-$\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$ $=\frac{\partial}{\partial\mathbf{W}}\left(\mathbf{X}^{(0)}+\sum_{k_{0}=1}^{K}(-1)^{k_{0}}\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)}\mathbf{W}^{k_{0}}\right)$
-$=\sum_{k_{0}=1}^{K}(-1)^{k_{0}}k_{0}\frac{\partial}{\partial\mathbf{W}^{k_{0}}}(\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)}\mathbf{W}^{k_{0}})\mathbf{W}^{k_{0}-1}\frac{\partial\mathbf{W}}{\partial\mathbf{W}}.$ (19)
-Finally, $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ can be derived as the multiplication of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-1)}}$ and $\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$:
-$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\sum_{k_{0}=1}^{K}\Bigg{[}(-1)^{k_{0}}k_{0}(\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)})^{\top}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)\mathbf{W}^{k_{0}-1}$
-$+\sum_{\ell=2}^{L_{R}-1}\binom{L_{R}-1}{\ell}\sum_{(k_{1},k_{2},\cdots,k_{\ell-1})}(-1)^{\sum_{i=1}^{\ell-1}k_{i}}k_{0}(\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)})^{\top}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}\mathbf{W}^{k_{0}-1}\Bigg{]}.$ (20)
-By the same procedure in the forward process, the time complexity of the first term in brackets is $O(k_{0}(|\mathcal{E}|d+d^{3})+|\mathcal{V}|d^{2})$.
-In the case of the second term, there are up to $2\ell K$ possible outcomes for $\sum_{i=1}^{\ell}k_{i}$.
-So the time complexity is $O(L_{R}+K\ell|\mathcal{V}|d^{2}+\ell^{2}K^{2}(d^{3}+|\mathcal{E}|d))$.
-So the overall time complexity of reverse process is $O((L_{D}^{2}+M^{2}L_{R}^{3})|\mathcal{E}|d+(L_{D}^{2}+ML_{R}^{2})|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$.
+$=$ $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{\ell=1}^{L_{R}-1}\binom{L_{R}-1}{\ell}\sum_{(m_{1},m_{2},\cdots,m_{\ell})}(-1)^{\sum_{i=1}^{\ell}m_{i}}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}\;,$ (18)
+where $m_{i}=1,\cdots,M$ for all $i$.
+To demonstrate Equation 18, we show part of the sequential multiplication process.
+Multiplying $\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}$ to the right side of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}$, we get:
+$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial}{\partial\mathbf{X}^{(-L_{R}+1)}}\left(\mathbf{X}^{(-L_{R}+1)}+\sum_{m=1}^{M}(-1)^{m}\hat{\mathbf{A}}^{m}\mathbf{X}^{(-L_{R}+1)}\mathbf{W}^{m}\right)$
+$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{m=1}^{M}(-1)^{m}\left(\hat{\mathbf{A}}^{m}\right)^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\left(\mathbf{W}^{m}\right)^{\top}\;,$ (19)
+which can also be obtained by substitute $L_{R}=2$ in Equation 18.
+Next, we multiply $\frac{\partial\mathbf{X}^{(-L_{R}+1)}}{\partial\mathbf{X}^{(-L_{R}+2)}}$ on the right side of Equation 19, yielding:
+$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}\frac{\partial\mathbf{X}^{(-L_{R}+1)}}{\partial\mathbf{X}^{(-L_{R}+2)}}$
+$=$ $\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{m=1}^{M}(-1)^{m}\left(\hat{\mathbf{A}}^{m}\right)^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\left(\mathbf{W}^{m}\right)^{\top}\right)\frac{\partial}{\partial\mathbf{X}^{(-L_{R}+2)}}\left(\mathbf{X}^{(-L_{R}+2)}+\sum_{m=1}^{M}(-1)^{m}\hat{\mathbf{A}}^{m}\mathbf{X}^{(-L_{R}+2)}\mathbf{W}^{m}\right)$
+$=$ $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{\ell=1}^{2}\binom{2}{\ell}\sum_{(m_{1},m_{2},\cdots,m_{\ell})}(-1)^{\sum_{i=1}^{\ell}m_{i}}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}.$ (20)
+Repeating the process above, we can derive Equation 18.
+Finally, $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ is derived by multiplying Equation 18 with $\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$:
+$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\sum_{m_{0}=1}^{M}(-1)^{m_{0}}m_{0}(\hat{\mathbf{A}}^{m_{0}}\mathbf{X}^{(0)})^{\top}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)\mathbf{W}^{m_{0}-1}$
+$+\sum_{\ell=1}^{L_{R}-1}\binom{L_{R}-1}{\ell}\sum_{(m_{0},m_{1},\cdots,m_{\ell})}(-1)^{\sum_{i=1}^{\ell}m_{i}}m_{0}(\hat{\mathbf{A}}^{m_{0}}\mathbf{X}^{(0)})^{\top}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}\mathbf{W}^{m_{0}-1}\;,$ (21)
+where $m_{i}=1,\cdots,M$ for all $i$.
+The time complexity of the first term is $O(M^{2}|\mathcal{E}|d+M|\mathcal{V}|d^{2}+M^{2}d^{3})$.
+In the case of the second term, the time complexity is $O(M^{2}L_{R}^{3}|\mathcal{E}|d+ML_{R}^{2}|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$, since there are up to $2\ell M$ possible outcomes for $\sum_{i=1}^{\ell}m_{i}$.
+Therefore, the overall time complexity of the reverse process is $O((L_{F}^{2}+M^{2}L_{R}^{3})|\mathcal{E}|d+(L_{F}^{2}+ML_{R}^{2})|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$.
 The memory complexity is $O(|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
@@
 $h(\mathbf{X}^{(\ell)})=\sigma(\frac{1}{K}\sum_{k=1}^{K}\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}\mathbf{W}^{(k)}).\;$ (23)
-In this case, the upper bound of the Lipschitz constant of $h$ is calculated as follows:
-$\text{Lip}(h)$ $=\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\frac{1}{K}\sum_{k=1}^{K}\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}\mathbf{W}^{(k)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}$ (23)
-$\leq\frac{1}{K}\sum_{k=1}^{K}\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}\lVert{\mathbf{W}^{(k)}}\rVert_{F}$ (24)
+In this case, the Lipschitz constant of $h$, $\text{Lip}(h)<1$ is satisfied if
+$\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\frac{1}{K}\sum_{k=1}^{K}\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}\mathbf{W}^{(k)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}<1\;.$ (24)
+The upper bound of left side is computed by:
+$\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\frac{1}{K}\sum_{k=1}^{K}\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}\mathbf{W}^{(k)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}$ $\leq\frac{1}{K}\sum_{k=1}^{K}\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}\lVert{\mathbf{W}^{(k)}}\rVert_{F}$
 $\leq\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F},\;$ (25)